### v1

# LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology

Abstract
Large language models (LLMs) and emerging agentic frameworks are beginning to transform single-cell biology by enabling natural-language reasoning, generative annotation, and multimodal data integration. However, progress remains fragmented across data modalities, architectures, and evaluation standards. LLM4Cell presents the first unified survey of 58 foundation and agentic models developed for single-cell research, spanning RNA, ATAC, multi-omic, and spatial modalities. We categorize these methods into five families—foundation, text-bridge, spatial/multimodal, epigenomic, and agentic—and map them to eight key analytical tasks including annotation, trajectory and perturbation modeling, and drug-response prediction. Drawing on over 40 public datasets, we analyze benchmark suitability, data diversity, and ethical or scalability constraints, and evaluate models across 10 domain dimensions covering biological grounding, multi-omics alignment, fairness, privacy, and explainability. By linking datasets, models, and evaluation domains, LLM4Cell provides the first integrated view of language-driven single-cell intelligence and outlines open challenges in interpretability, standardization, and trustworthy model development.
LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology
Sajib Acharjee Dipthanks:  Equal contribution.1, Adrika Zafor11footnotemark: 12, Bikash Kumar Paul1 Uddip Acharjee Shuvo3, Muhit Islam Emon1, Xuan Wang1, Liqing Zhangthanks:  Correspondence: sajibacharjeedip@vt.edu, lqzhang@cs.vt.edu.1,4 1Department of Computer Science, Virginia Tech, Blacksburg, VA, USA 2Department of Computational Modeling and Data Analytics, Virginia Tech, Blacksburg, VA, USA 3Institute of Information and Technology, University of Dhaka, Dhaka, Bangladesh 4Fralin Biomedical Research Institute at VTC: Cancer Research Center, Washington DC, USA
1 Introduction
Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text. In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework. Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis. Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
Despite this progress, the field remains fragmented. Models differ widely in data modality, supervision type, and evaluation standard—ranging from scRNA-seq to ATAC, multi-omic, and spatial measurements—hindering cross-model comparison and reproducibility. Benchmarking is inconsistent, datasets are unevenly distributed across modalities, and agentic frameworks lack standardized ways to measure reasoning correctness or biological grounding. Without a systematic synthesis, it is difficult to assess what current single-cell LLMs truly understand or where they fall short.
To address these challenges, we introduce LLM4Cell, a comprehensive survey of large language and agentic models for single-cell biology. We analyze 58 representative methods spanning five methodological families—Foundation, Text-Bridge, Spatial/Multimodal, Epigenomic, and Agentic—and organize them across eight major tasks, from annotation and trajectory inference to perturbation and drug-response modeling. Complementing this, we compile over 40 publicly available datasets across RNA, ATAC, multi-omic, spatial, perturbation, and plant domains, and evaluate each model using a ten-dimension rubric covering grounding, fairness, scalability, and interpretability.
Contributions.
LLM4Cell provides: (1) a modality-balanced registry of $\sim$40 benchmark datasets for large-model training and evaluation; (2) a unified taxonomy of 58 foundation and agentic models across eight analytical tasks; (3) a ten-dimension domain rubric capturing biological grounding, generalization, and ethics; and (4) a critical discussion of open challenges in cross-modal alignment, reasoning, and trustworthy AI for cell biology.
2 Related Work
The intersection of large language models (LLMs) and single-cell biology has gained rapid momentum, motivating several early surveys and benchmarks. The ACL 2025 survey on foundation models for single-cell biology (Zhang et al., 2025a) outlines pretrained and fine-tuned architectures across protein, gene, and cell representations. Broader reviews such as Lan et al. (2025) and Bian et al. (2024b) discuss “large cellular models” (LCMs) and trends in scaling, tokenization, and transfer learning across biological data. Benchmark studies including Single-Cell Omics Arena (Liu et al., 2024) and CellVerse (Zhang et al., 2025b) evaluate LLMs on annotation and question answering, while works like scInterpreter (Guo et al., 2023) and scExtract (Wu and Tang, 2025) demonstrate practical integration into annotation and curation pipelines.
However, existing reviews remain limited in scope. They focus on architectures or prompting but give little attention to the datasets underpinning model training and evaluation—especially those covering spatial, perturbation, or multimodal modalities. Standardized benchmarks, domain-aware metrics, and reproducible splits are seldom addressed, and plant single-cell datasets—key for cross-kingdom generalization—are almost entirely absent. Moreover, agentic reasoning and tool-augmented systems are typically treated as isolated examples rather than a coherent methodological class, leaving the relationship among data modalities, modeling paradigms, and biological domains fragmented.
LLM4Cell advances this literature by providing an integrated, data-centric synthesis linking models, datasets, and evaluation domains. Unlike prior descriptive taxonomies, it unifies over forty public datasets and fifty-eight models across modalities and tasks, assessing their biological grounding, fairness, and scalability. In doing so, LLM4Cell establishes a reproducible foundation for cross-modal benchmarking and future research on trustworthy, autonomous single-cell intelligence.
3 Datasets
Progress in large language models for single-cell biology relies on the rapid growth of high-quality datasets across transcriptomic, epigenomic, multimodal, and spatial domains. Our survey compiles over forty public resources spanning five major modalities—RNA, ATAC, multi-omic, spatial, and perturbation/drug-response—plus emerging plant single-cell atlases. These datasets underpin model pretraining, evaluation, and cross-modal reasoning for annotation, integration, trajectory inference, perturbation modeling, and spatial mapping.
Transcriptomic atlases such as Tabula Sapiens, Tabula Muris  (Consortium* et al., 2022), and the Human Cell Atlas  (Travaglini et al., 2020) remain dominant for foundation-model training and ontology-aware annotation. Chromatin-accessibility data (e.g. Cusanovich mouse atlas (Cusanovich et al., 2018), human adult/fetal scATAC (Domcke et al., 2020)) map regulatory states but remain sparse and heterogeneous. Multi-omic resources (e.g. TEA-seq  (Swanson et al., 2021), DOGMA-seq  (Mimitou et al., 2021), CITE-seq (Stoeckius et al., 2017)) link RNA, ATAC, and protein modalities, providing supervision for cross-view alignment. Spatial technologies (e.g. Visium  (Oliveira et al., 2025), Slide-seqV2  (Stickels et al., 2021), MERFISH  (Chen et al., 2015), Stereo-seq  (Chen et al., 2022)) connect molecular profiles to tissue architecture, while functional datasets (e.g. Perturb-seq (Replogle et al., 2022)) benchmark causal reasoning. Plant resources (e.g. scPlantDB  He et al. (2024), Arabidopsis E-CURD-4  (Shulse et al., 2019)) extend modeling beyond animal systems, opening cross-kingdom evaluation.
Despite this diversity, major gaps persist. (i) RNA atlases vastly outscale other modalities. (ii) Benchmark fragmentation and inconsistent metadata hinder reproducibility. (iii) Privacy constraints limit access to clinical spatial datasets. (iv) Non-human and plant data remain scarce. (v) Paired and tri-modal resources provide ideal testbeds for next-generation multimodal and agentic LLMs. Table LABEL:tab:datasets-rna, LABEL:tab:datasets-atac, LABEL:tab:datasets-multiome, LABEL:tab:datasets-spatial, LABEL:tab:datasets-perturb, LABEL:tab:datasets-plant (Appendix D) list representative datasets with tasks, scale, and source links for reproducible benchmarking.
4 Model Taxonomy
Foundation Models Self-supervised (masked / next-token) Multi-task / adapters Cross-species scGPT, Geneformer, scFoundation, scBERT, tGPT, scTrans, scRDiT, scGraphformer, scGFT CellFM, scPRINT, scMulan, scMoFormer, scmFormer iSEEEK, UCE, GeneCompass Text-Bridge & Ontology LLMs Cell-to-text Text-conditioned embeddings Ontology / prompt alignment Cell2Sentence (C2S), C2S-Scale, Cell2Text GenePT (GENEPT), scELMo, scInterpreter CellLM, QuST-LLM Spatial & Multimodal FMs RNA + Spatial RNA + ATAC + Protein Histology-conditioned TransformerST, stFormer, scGPT-spatial, Spatial2Sentence, spaCCC scMMGPT, OmiCLIP, FmH2ST spaLLM, HEIST, Nicheformer Epigenomic & Regulatory FMs Chromatin accessibility Efficient / long-range Cross-omic regulatory EpiFoundation, EpiBERT, EpiAttend GeneMamba, scMamba ChromFound, GET Agentic & Reasoning Frameworks LLMs as annotation agents Reasoning + tools Interactive / multi-agent scAgent, LICT EpiAgent, ChatCell, scExtract CellVerse, Teddy, Pilot
Figure 1: Hierarchical taxonomy for LLM4Cell. Color-coded families expand into sub-branches and representative models, tracing the progression from foundation pretraining to multimodal and agentic reasoning frameworks. References are omitted for visibility and included in the appendix method comparison table.
Large language models are rapidly reshaping single-cell biology, producing diverse architectures, pretraining schemes, and reasoning frameworks. We categorize 58 representative methods into five methodological families—Foundation Models, Text-Bridge LLMs, Spatial and Multimodal Models, Epigenomic Models, and Agentic Frameworks—based on their core design and data modality. These span the progression from gene-level embeddings to multimodal and autonomous systems capable of biological reasoning.
Our taxonomy is organized along five orthogonal dimensions: (i) Modality (RNA, ATAC, multi-omic, spatial, or hybrid), (ii) Grounding type (atlas, ontology, or marker-based), (iii) Agentic capability (multi-step reasoning or autonomous orchestration), (iv) Primary task (annotation, trajectory, perturbation, integration, etc.), and (v) Domain quality, represented by ten quantitative scores for grounding, fairness, scalability, and interpretability.
Figure 1 summarizes this taxonomy, with detailed attributes in Appendix Table LABEL:tab:methods-full. Foundation models currently dominate in scope and adoption, while emerging spatial, epigenomic, and agentic frameworks signal a shift toward contextual, reasoning-driven single-cell intelligence.
4.1 Foundation Models
Foundation models learn transferable cell and gene embeddings directly from large-scale scRNA-seq without explicit labels. Representative systems such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) pretrain on multi-tissue atlases exceeding one million cells, using masked-gene or rank-based reconstruction to capture expression context. Variants like tGPT  (Shen et al., 2023), scBERT  (Yang et al., 2022), and scGraphformer  (Fan et al., 2024) treat genes as tokens or graph nodes, while cross-species models (UCE  (Rosen et al., 2023), GeneCompass (Yang et al., 2024b)) apply contrastive alignment across homologous genes. These models underpin most single-cell LLM pipelines, offering strong transfer for annotation and integration but limited ontology grounding and explainability, motivating later text-bridged and reasoning frameworks.
Figure 2: Task and domain distribution across spatial transcriptomics models. Most emphasize biological grounding, scalability, and annotation tasks.
4.2 Text-Bridge LLMs
Text-Bridge models couple molecular embeddings with biomedical language to ground single-cell representations in semantics and ontology. Systems such as scELMo  (Liu et al., 2023), CellLM  (Zhao et al., 2023), and GenePT  (Chen and Zou, 2024) align gene or cell embeddings with textual descriptors, while Cell2Sentence  (Levine et al., 2024) and Cell2Text  (Kharouiche et al., 2025) translate expression profiles into natural-language summaries. Using dual-encoder or encoder–decoder architectures, these frameworks employ contrastive or prompt-based alignment between molecular and text encoders (e.g., BioBERT). They enhance interpretability and enable zero-shot annotation but depend heavily on curated corpora and remain non-agentic, serving as a bridge between foundation and reasoning frameworks.
4.3 Spatial and Multimodal Models
Spatial and multimodal frameworks integrate gene expression with spatial coordinates, histology, or additional omics to capture tissue architecture. Representative models include TransformerST  (Lu et al., 2024), spaLLM (Ji et al., 2024), and OmiCLIP (Cui et al., 2025) for spatial mapping, and scMMGPT (Shi et al., 2025) and FmH2ST for RNA–ATAC–protein integration. They use multi-branch Transformers with modality-specific encoders and cross-attention fusion, trained on datasets such as Visium DLPFC and MERFISH. These models achieve strong spatial alignment and biological realism but face heterogeneous resolutions, limited open benchmarks, and high computational cost.
4.4 Epigenomic Models
Epigenomic foundation models extend LLM concepts to chromatin-accessibility and regulatory data such as scATAC-seq. EpiFoundation  (Wu et al., 2025), EpiBERT (Javed et al., 2025), and EpiAttend (Li et al., 2022) learn cis-regulatory patterns from ENCODE and other compendia, while GeneMamba and scMamba use efficient state-space layers for long-range dependency modeling. Cross-omic variants (ChromFound (Jiao et al., 2025), GET (Fu et al., 2025)) jointly embed RNA and ATAC to infer gene-regulatory networks. These models improve biological grounding but remain constrained by sparse data and lack unified benchmarks across regulatory modalities.
4.5 Agentic Frameworks
Agentic systems integrate pretrained models with reasoning modules for autonomous single-cell analysis. Frameworks such as scAgent (Mao et al., 2025), CellVerse (Zhang et al., 2025b), and Teddy (Chevalier et al., 2025) combine domain-specific encoders with LLM controllers that plan tasks, query ontologies, and interface with tools or APIs. EpiAgent (Chen et al., 2025b) extends this paradigm to regulatory genomics. These systems enable dialogue-based annotation and multi-step reasoning but lack standardized benchmarks for reasoning fidelity and rely on underlying LLM reliability.
5 Task-specific Applications
While the previous section categorized models by architectural family, here we analyze them through the lens of biological objectives shown in Figure 2 and Appendix Figure 3. Across the 58 methods surveyed, we identify eight recurring tasks that define single-cell modeling pipelines: (1) annotation and ontology mapping, (2) trajectory and perturbation modeling, (3) multi-omic integration, (4) spatial mapping and deconvolution, (5) regulatory-network and pathway inference, (6) cross-species translation, (7) generative simulation, and (8) drug-response prediction. Most models span multiple tasks-for example, scGPT and Geneformer support both annotation and perturbation prediction-reflecting the convergence between representation learning and functional reasoning.
This task-centric view complements the architectural taxonomy by revealing how foundation and agentic systems differ in operational scope. Foundation models dominate annotation, integration, and trajectory inference, whereas emerging text-bridge and agentic systems extend to knowledge-grounded reasoning and dynamic planning. Spatial and epigenomic frameworks contribute specialized capabilities in tissue mapping and chromatin-level interpretation.
5.1 Annotation and Ontology Mapping
Annotation is the most common single-cell task, spanning automated cell-type labeling, ontology alignment, and cross-dataset harmonization. Foundation models such as scGPT, scFoundation, and scBERT achieve high accuracy on atlases like Tabula Sapiens and the Human Cell Atlas. Text-bridge models (CellLM  (Zhao et al., 2023), scELMo  (Liu et al., 2023), GenePT  (Chen and Zou, 2024)) add ontology-based interpretability, and agentic systems (scAgent) perform multi-step annotation reasoning via LLM controllers. Models are fine-tuned on curated references (e.g., Azimuth PBMC, HCA Lung) using masked-gene or contrastive objectives. Evaluation uses label accuracy or ARI; key limitations include rare-cell detection, cross-species consistency, and lack of reasoning benchmarks.
5.2 Trajectory and Perturbation Modeling
Trajectory modeling captures dynamic state transitions and causal responses to interventions. Geneformer  (Theodoris et al., 2023), scGPT (Cui et al., 2024), and scRDiT  (Dong et al., 2025) model temporal or perturbation trajectories using denoising and contrastive objectives, trained on datasets such as Replogle 2022 Perturb-seq and sci-Plex. Epigenomic models (EpiFoundation (Wu et al., 2025) ,EpiAgent (Chen et al., 2025b)) extend to regulatory responses through chromatin context and ontology reasoning. Metrics include expression-profile correlation and target-gene recovery. While these models predict single perturbations effectively, performance degrades for combinatorial or long-range effects, underscoring the need for unified causal benchmarks.
5.3 Multi-omic Integration
Multi-omic integration seeks unified representations across paired RNA, ATAC, and protein modalities to capture gene-regulatory and signaling relationships. Representative models include scMMGPT (Shi et al., 2025), GET (General Expression Transformer) (Fu et al., 2025), ChromFound (Jiao et al., 2025), and epigenomic-aware variants such as EpiFoundation (Wu et al., 2025) and EpiBERT (Javed et al., 2025). These systems train on joint datasets like 10x Multiome PBMC 10k, TEA-seq, DOGMA-seq, and ASAP-seq, using cross-modal transformers or contrastive alignment between expression and accessibility features.
Most architectures employ modality-specific encoders with shared latent spaces or token-type embeddings; scMGPT applies masked-reconstruction across modalities, while GET learns cross-attention between gene and chromatin tokens. Compared with unimodal models, they show improved cell-type alignment and batch correction but face data-scale imbalance and sparse modality overlap. Evaluation commonly uses modality-matching accuracy and latent-space correlation. Despite strong integration performance, interpretability and standard benchmarks remain limited, and scalability to tri-modal data is an active challenge.
5.4 Spatial Mapping and Deconvolution
Spatial mapping links molecular profiles to their tissue locations, enabling inference of cell organization and microenvironmental context. Representative models include TransformerST  (Lu et al., 2024), spaLLM , OmiCLIP, FmH2ST, HEIST  (Madhu et al., 2025), and Spatial2Sentence (Chen et al., 2025a), trained on datasets such as Visium DLPFC, Slide-seqV2, MERFISH, and Xenium. They combine spatial coordinates, histology features, and gene-expression embeddings through cross-attention or contrastive alignment to reconstruct cell or spot-level maps.
TransformerST performs axial attention between gene and spatial axes for spot-to-cell deconvolution, while spaLLM and OmiCLIP align histology and transcriptomic features using language-guided contrastive learning. Models like HEIST and FmH2ST fuse H&E images with RNA profiles for tissue-domain segmentation. Evaluation typically measures spatial correlation, clustering accuracy, or cell-type F1 against curated annotations. Although these models improve contextual understanding of tissue architecture, limited spatial benchmarks and heterogeneous resolutions constrain systematic comparison across methods.
5.5 Regulatory Network and Pathway Inference
Regulatory-network and pathway inference focuses on identifying gene–gene or enhancer–promoter dependencies that drive transcriptional programs. Representative models include GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025), EpiFoundation (Wu et al., 2025), EpiBERT (Javed et al., 2025), ChromFound (Jiao et al., 2025), and GET (Fu et al., 2025), trained on large-scale ATAC and multiome atlases such as ENCODE, human cCRE datasets, and 10x Multiome PBMC. These models treat chromatin regions or genes as tokens and learn context-dependent regulatory relationships through self- or cross-attention.
GeneMamba (Qi et al., 2025a) and scMamba employ state-space layers for efficient modeling of long-range chromatin dependencies, while EpiBERT and EpiFoundation use masked-region reconstruction to capture enhancer–promoter coupling. ChromFound and GET link RNA and accessibility signals to infer transcription-factor activity and causal directionality. Evaluation typically uses motif enrichment, AUROC for known interactions, or pathway-level overlap with KEGG or Reactome references. These models improve biological grounding and causal interpretability but remain constrained by sparse training data and lack of unified ground-truth regulatory benchmarks.
5.6 Cross-Species Translation
Cross-species translation aims to transfer cell and gene representations between organisms, enabling comparative and evolutionary analysis of single-cell atlases. Representative models include iSEEEK  (Shen et al., 2022), UCE (Universal Cell Embeddings)  (Rosen et al., 2023), GeneCompass  (Yang et al., 2024b), and scPlantLLM  (Cao et al., 2025), trained on paired or homolog-mapped datasets such as tthe Mouse Cell Atlas, Tabula Muris, and scPlantDB. These systems learn alignment between orthologous genes or conserved cell states to support zero-shot annotation across species.
iSEEEK  (Shen et al., 2022) integrates over 11 million human and mouse cells using shared gene-token vocabularies, while UCE  (Shen et al., 2023) applies masked autoencoding and contrastive alignment across ortholog groups. GeneCompass  (Yang et al., 2024b) extends this with knowledge-informed pretraining that links species via ontology-derived embeddings, and scPlantLLM adapts the paradigm to plant single-cell data across 17 species. Evaluation uses transfer accuracy and embedding consistency across species pairs. Despite improved cross-domain generalization, these models remain limited by incomplete ortholog mappings and domain shifts in sequencing depth and tissue composition.
5.7 Generation and Simulation
Generative modeling aims to synthesize realistic single-cell profiles, simulate perturbations, or reconstruct missing modalities. Representative approaches include scGPT, scFoundation, CellFM  (Zeng et al., 2025), and Geneformer, which employ decoder or diffusion-style architectures to model the probability distribution of gene-expression states. These models are pretrained on large atlases such as Tabula Sapiens and HCA, then fine-tuned to generate new cell embeddings or perturbation outcomes.
scGPT and scFoundation use masked-token prediction extended with generative decoding to simulate unseen cells or conditions, while CellFM introduces multimodal conditioning for cross-tissue synthesis. Generative evaluation typically measures distributional similarity (KL divergence, FID-like embedding distance) or recovery of known cell-type proportions. Such models enable scalable in-silico experiments and data augmentation but face challenges in controlling biological realism and avoiding overfitting to dominant cell types.
5.8 Drug-Response Prediction
Drug-response prediction models aim to infer cellular transcriptional outcomes following chemical or genetic perturbations, supporting virtual screening and mechanism discovery. Representative frameworks include Geneformer (Theodoris et al., 2023), scGPT  (Cui et al., 2024), EpiFoundation (Wu et al., 2025), and EpiAgent (Chen et al., 2025b), trained or evaluated on large-scale pharmacotranscriptomic resources such as sci-Plex, Replogle 2022 Perturb-seq, and the ARC Virtual Cell Challenge. These systems learn conditional embeddings that map control to perturbed states, enabling zero-shot prediction of unseen compounds or targets.
Geneformer predicts post-treatment expression profiles using masked-token denoising, while scGPT and EpiFoundation integrate modality-specific conditioning to capture epigenetic or pathway context. EpiAgent further incorporates ontology-guided reasoning to link molecular features with literature-derived drug–gene relationships. Evaluation relies on Pearson correlation or top-$k$ recovery of known targets between predicted and observed responses. Although these models demonstrate strong predictive capacity for single-agent perturbations, performance degrades for combinatorial treatments and unseen pathways, highlighting the need for richer multimodal pharmacogenomic benchmarks.
Cross-task observations. Across all eight task domains, clear convergence patterns emerge. Foundation models dominate annotation, integration, and generative simulation owing to large-scale pretraining on transcriptomic atlases, whereas text-bridge systems extend interpretability through ontology grounding and natural-language supervision. Spatial and epigenomic models expand biological realism by capturing tissue architecture and regulatory logic, while agentic frameworks introduce multi-step reasoning and tool orchestration that connect these modalities into autonomous workflows. Despite methodological diversity, evaluation remains fragmented-annotation and integration enjoy standardized benchmarks, but trajectory, drug-response, and agentic reasoning lack common protocols. Cross-species and privacy-aware tasks are the least represented, reflecting data scarcity and regulatory constraints. Overall, task analysis reveals a rapidly maturing ecosystem in which foundation and multimodal pretraining supply general representations, and emerging agentic paradigms begin to operationalize biological reasoning. The following section quantifies these trends through ten domain-level dimensions encompassing biological grounding, fairness, scalability, and interpretability.
6 Domain Analysis
We evaluate fifty-eight methods across ten domain dimensions capturing reliability, generalization, and interpretability. Each score reflects published evidence of performance or explicit design alignment, summarized in Appendix Table LABEL:tab:methods-full and Figure 2. Together, these dimensions provide a composite view of model maturity beyond task accuracy.
6.1 Biological Grounding
Most foundation and multimodal models achieve strong biological grounding by encoding marker genes, pathways, or regulatory regions. scGPT, Geneformer, and EpiFoundation leverage atlas-scale training to recover known cell types and pathway enrichments. Spatial frameworks such as TransformerST and OmiCLIP add tissue-context grounding via histology alignment. However, textual and ontology grounding remain inconsistent, and biological semantics are rarely explicit in model objectives.
6.2 Batch Effects and Heterogeneity
Robustness to batch effects varies widely. scFoundation  (Hao et al., 2024), CellFM  (Zeng et al., 2025), and iSEEEK  (Shen et al., 2022) mitigate donor and protocol variation through large-scale multi-domain pretraining, while others rely on explicit batch tokens or adversarial alignment. True cross-platform generalization remains limited; even large models show degradation when tested on unseen sequencing chemistries or tissue types.
6.3 Multi-omics Alignment
Models such as scMGPT  (Palayew et al., 2025), GET  (Fu et al., 2025), and ChromFound  (Jiao et al., 2025) demonstrate consistent alignment across RNA, ATAC, and protein features using cross-attention or contrastive fusion. Spatial hybrids (HEIST, FmH2ST) further connect molecular and visual domains. While alignment quality improves, limited paired datasets and uneven modality coverage constrain reproducibility and benchmarking.
6.4 Trajectory and Perturbation
Dynamic inference is strongest in generative transformers (Geneformer, scGPT) and epigenomic variants (EpiAgent). They predict post-perturbation expression or accessibility shifts with reasonable fidelity but struggle to model multi-target or time-continuous responses. Few methods quantify causal uncertainty or biological plausibility.
6.5 Cross-Species and Cross-Tissue Generalization
Cross-organism transfer remains challenging. iSEEEK  (Shen et al., 2022), UCE  (Rosen et al., 2023), and GeneCompass  (Yang et al., 2024b) align human and mouse cell embeddings via ortholog mapping, while scPlantLLM  (Cao et al., 2025) extends this to plants. Performance drops sharply for divergent taxa, reflecting incomplete ortholog tables and bias toward human datasets.
6.6 Atlas Fairness and Representation Balance
Dataset imbalance affects nearly all models. Human and immune-cell–dominant atlases overrepresent certain tissues and demographics. Teddy (Chevalier et al., 2025) introduces benchmarking for fairness and representation balance, but fairness metrics are rarely adopted. No model yet enforces demographic or tissue-specific parity during training.
6.7 Explainability and Interpretability
Text-bridge systems (GenePT, CellLM, Cell2Text  (Kharouiche et al., 2025)) improve interpretability by linking embeddings to ontology terms. Agentic systems (scAgent) extend this via natural-language reasoning and tool explanations. Most foundation models still rely on attention visualization, providing limited causal transparency.
6.8 Privacy and Ethics
Data privacy remains underexplored. Synthetic generation (scGFT  (Nouri, 2025)) and open-access chat systems (ChatCell  (Fang et al., 2024)) raise concerns over re-identification and data leakage. No single-cell LLM currently implements federated or privacy-preserving learning, and ethical guidance for multi-omic sharing remains informal.
6.9 Scalability and Efficiency
Scaling trends mirror NLP: efficient transformers (xTrimoGene (Gong et al., 2023)) and state-space models (GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025)) reduce memory cost while retaining accuracy. CellFM  (Zeng et al., 2025) demonstrates training across 100 M cells, showing feasibility of atlas-scale learning. Still, high compute requirements limit accessibility for most research groups.
6.10 Emerging Paradigms and Agentic Behavior
Recent frameworks (scAgent  (Mao et al., 2025), CellVerse  (Zhang et al., 2025b), EpiAgent (Chen et al., 2025b)) introduce reasoning and autonomous decision pipelines, integrating LLM controllers with specialized encoders. These systems achieve the highest scores in explainability and cross-modal planning but lack standardized evaluation of reasoning fidelity or reproducibility.
7 Open Problems
7.1 Trust and Validation
Evaluation across modalities remains inconsistent. Metrics often emphasize reconstruction over biological plausibility, and few studies offer independent replication. Community-curated leaderboards and standardized validation sets are essential for credible comparison.
7.2 Data and Bias
Training corpora are dominated by human and mouse atlases, limiting cross-species and clinical generalization. Rare-cell, plant, and microbial systems remain underrepresented, reinforcing demographic and biological bias. Diverse, balanced datasets are critical for equitable model development.
7.3 Cross-Modal and Dynamic Modeling
True integration of RNA, ATAC, spatial, and temporal modalities is still elusive. Existing models handle only pairwise combinations, while four-way fusion remains computationally infeasible. Unified tokenization and multimodal pretraining pipelines could enable consistent cross-domain reasoning.
7.4 Interpretability and Causality
LLM embeddings capture statistical correlation but rarely mechanistic insight. Bridging attention to regulatory or causal networks requires explicit reasoning layers and experimental grounding. Combining symbolic or causal discovery modules with transformers is a key next step.
7.5 Ethics and Privacy
Open single-cell data sharing raises privacy and dual-use concerns. Few models include audit trails or consent-aware governance. Adapting federated and differential-privacy methods to biological data is urgently needed.
7.6 Agentic and Interactive Systems
Agentic frameworks can plan and reason but lack reliable benchmarks. Tasks such as multi-step reasoning, dataset retrieval, and model orchestration remain fragile. Robust metrics for reasoning accuracy, safety, and reproducibility will be crucial for trustworthy biological AI.
8 Conclusions
LLM4Cell presents the first unified survey of large language and agentic models for single-cell biology, linking datasets, architectures, and evaluation domains within a common framework. By analyzing 58 models and over 40 public datasets across RNA, ATAC, multi-omic, spatial, and perturbation modalities, we show how foundation-scale pretraining, multimodal grounding, and agentic reasoning are reshaping single-cell analysis. Our taxonomy and evaluation rubric reveal a transition from purely statistical modeling toward language-driven, interpretable, and increasingly autonomous systems that connect molecular data with biological knowledge. We hope this synthesis provides a reproducible reference for benchmarking, model selection, and the design of next-generation cellular foundation and reasoning models.
Limitations
Despite its breadth, this study has several limitations. First, reported performance metrics vary widely across publications, preventing consistent quantitative comparison. Second, our ten-dimension rubric captures qualitative trends but not standardized numerical rankings. Third, access restrictions and licensing prevented inclusion of certain clinical or proprietary spatial datasets, while non-animal single-cell resources remain scarce. Fourth, model efficiency, compute cost, and hyperparameter sensitivity were not systematically analyzed. Finally, as multimodal and reasoning frameworks evolve rapidly, LLM4Cell reflects a snapshot in time rather than a static benchmark. Future work should establish community leaderboards, causal interpretability tests, and reasoning benchmarks to advance trustworthy and reproducible single-cell intelligence.
Appendix A Data Collection
A.1 Search Strategy
To construct a comprehensive registry of large language models (LLMs) and agentic frameworks in single-cell biology, we systematically screened both peer-reviewed and preprint literature from 2020–2025 across multiple sources. Searches were conducted on PubMed, Google Scholar, arXiv, and Semantic Scholar using combinations of domain- and method-specific keywords, including:
  • General: “large language model”, “foundation model”, “LLM”, “multimodal model”, “generative AI”
  • Domain-specific: “single-cell”, “scRNA-seq”, “scATAC”, “multiome”, “spatial transcriptomics”, “perturb-seq”, “cell atlas”, “biomedical foundation model”
  • Integration / reasoning: “ontology alignment”, “text-bridge”, “agentic framework”, “biological reasoning”, “cross-modal integration”, “annotation model”
After removing duplicates, the combined queries returned approximately 8,020 papers from 2020–2025. Of these, 5,510 remained after filtering for English-language articles with available abstracts. We manually reviewed titles, abstracts, and model descriptions to identify works directly involving large-scale pretrained models, multimodal transformers, or language-based reasoning systems applied to single-cell or cellular-level omics data.
A.2 Inclusion and Screening Criteria
Studies were included if they met at least one of the following criteria:
  1. 1.
    Introduced or evaluated a large pretrained model ($\geq$10M parameters) for single-cell or multi-omic analysis.
  2. 2.
    Employed textual grounding, ontology prompts, or natural-language interfaces to interpret single-cell data.
  3. 3.
    Proposed agentic or reasoning-based frameworks for biological tasks (annotation, trajectory, perturbation, or integration).
  4. 4.
    Released code, preprints, or benchmarks relevant to cellular foundation models.
Methodological, biological, and dataset-based duplicates were merged under representative entries. This curation yielded 58 distinct models across five methodological families: foundation, text-bridge, spatial/multimodal, epigenomic, and agentic frameworks. Each was annotated with task domain, modality, training data, and evaluation scope (Appendix Table LABEL:tab:methods-full).
A.3 Inclusion of Preprints
We included arXiv and bioRxiv preprints to ensure coverage of recent and influential models not yet published in peer-reviewed venues. In single-cell research, the majority of foundational architectures (e.g., scGPT, scFoundation, CellLM) initially appeared as preprints months before journal acceptance. Given the field’s rapid evolution and small number of LLM-scale models, preprints represent essential early contributions to reproducibility and transparency. Each preprint was manually cross-verified for active GitHub or Zenodo links to confirm technical validity and data availability.
A.4 Resulting Corpus
The final corpus spans 58 methods across 37 unique institutions, 40+ publicly accessible datasets, and 8 major analytical tasks. This dataset–method pairing forms the empirical foundation for the taxonomic and domain analyses presented in the main paper.
Appendix B Extended Dataset Summaries
B.1 RNA Atlases
Foundational scRNA-seq atlases such as Tabula Sapiens (>1 M cells, 24 tissues) and Tabula Muris Senis (mouse, multi-organ across aging) provide references for annotation and batch-effect evaluation. Organ-specific datasets like the Human Lung Cell Atlas (HLCA) and the Allen Brain Map enable domain-specific benchmarking and ontology-aware evaluation.
B.2 Chromatin and ATAC Data
The Cusanovich mouse sci-ATAC atlas, human adult/fetal scATAC atlases, and joint RNA–ATAC modalities (SHARE-seq, SNARE-seq2) characterize regulatory landscapes supporting trajectory and GRN inference, though sparsity limits large-scale pretraining.
B.3 Multiome and Tri-Modal Data
Datasets such as TEA-seq, DOGMA-seq, ASAP-seq, and CITE-seq pair transcriptomic, epigenomic, and proteomic modalities, enabling cross-modal translation and representation learning. The Multiome Benchmark Pack aggregates >40 datasets for standardization but remains smaller than large RNA atlases.
B.4 Spatial Transcriptomics and Imaging
Platforms including 10x Visium/HD, Slide-seqV2, and DBiT-seq provide spatial gene-expression maps, while imaging technologies (MERFISH, STARmap, Stereo-seq, CosMx, Xenium) capture near single-molecule resolution for tissue-level reasoning.
B.5 Perturbation and Drug-Response Screens
Large CRISPR-based datasets (Replogle 2022 Perturb-seq, Norman 2019, sci-Plex) and community benchmarks (Virtual Cell Challenge) quantify transcriptional responses for causal and policy evaluation in agentic frameworks.
B.6 Plant Single-Cell Datasets
scPlantDB (67 datasets, 17 species, 2.5 M cells) and PlantscRNAdb provide unified plant atlases; Arabidopsis E-CURD-4 and transgenic tobacco scRNA-seq datasets enable cross-kingdom and stress-response modeling.
B.7 Critical Observations
(1) RNA atlases dominate in scale; (2) Chromatin and spatial data are fragmented; (3) Privacy and licensing restrict clinical data; (4) Cross-species coverage is limited; (5) Paired modalities offer promising benchmarks for next-generation LLMs.
Representative datasets and links are provided in Tables A1–A3.
Appendix C Extended Details on Model Taxonomy
C.1 Foundation Models (Extended Details)
Foundation-scale models constitute the base layer for language-driven single-cell analysis. They are trained on molecular profiles—typically scRNA-seq or integrated multi-organ datasets—to learn generalizable representations of cellular states, gene programs, and perturbations.
Training corpora and scale.
Most use atlases containing $10^{6}$$10^{8}$ cells. scGPT and scFoundation were pretrained on human–mouse atlases from Tabula Sapiens and HCA; Geneformer aggregates thousands of GEO datasets; iSEEEK and CellFM span 11–17 organs and species, enabling cross-tissue generalization.
Architectural variations.
Most adopt Transformer backbones with gene tokens as contextual units. tGPT and scBERT encode ranked or discrete gene sequences, while Geneformer and scFoundation use masked-token modeling. Hybrid designs such as scGraphformer  (Fan et al., 2024) and scRDiT  (Dong et al., 2025) incorporate graph or diffusion-based attention, improving spatial and regulatory context.
Learning objectives.
Common objectives include masked-gene prediction, expression denoising, rank-based reconstruction, and cross-species contrastive learning (UCE, GeneCompass). Optimization typically minimizes cosine or KL divergence across gene embeddings, similar to masked-language modeling in NLP.
Applications and limitations.
Foundation embeddings transfer effectively to annotation, integration, and trajectory inference but lack explicit biological grounding. Ontology alignment and reasoning remain external, and interpretability is largely post-hoc (attention heatmaps, gene-ranking). These limitations prompted the development of Text-Bridge and Agentic frameworks discussed in later sections.
C.2 Text-Bridge LLMs (Extended Details)
These models explicitly connect molecular and textual modalities via ontology labels, pathway terms, or literature embeddings. scELMo fuses ELMo-based gene metadata with cell latent vectors; CellLM uses ontology-aware prompts for natural-language annotation; Cell2Sentence and Cell2Text map gene-expression vectors to descriptive sentences through cross-modal contrastive loss; and GenePT jointly pre-trains on PubMed to embed genes and diseases in the same space. Architectures typically adopt Transformer encoders for molecular data and domain-tuned BERT variants for text, trained with contrastive or cosine-similarity objectives. These systems rank highest in Explainability and Emerging paradigms scores (Appendix Table LABEL:tab:methods-full) but remain limited by vocabulary coverage and the lack of full multi-omic grounding.
C.3 Spatial and Multimodal Models (Extended Details)
Spatial models combine molecular and positional information to learn tissue-aware embeddings. TransformerST performs cross-scale attention between spatial spots and gene tokens; spaLLM aligns histology-derived captions with expression vectors using LLM guidance; and OmiCLIP links image and molecular embeddings via contrastive pre-training. Multi-omic extensions such as scMMGPT and FmH2ST integrate RNA, ATAC, and protein modalities within a shared latent space. Grounding typically relies on marker-based or atlas alignment rather than textual ontologies. Evaluation metrics include spot-level reconstruction accuracy and correlation with histological segmentation (DLPFC, Visium HD). These models show high biological grounding but moderate explainability due to visual-feature opacity.
C.4 Epigenomic Models (Extended Details)
Inputs include chromatin-accessibility matrices, motif sequences, and enhancer–promoter links. EpiFoundation pre-trains on the ENCODE scATAC compendium; EpiBERT encodes open-chromatin sequences using masked-region objectives; and EpiAttend models enhancer–promoter coupling through cross-region attention. State-space architectures (scMamba, GeneMamba) enable efficient context propagation over tens of thousands of genomic peaks. ChromFound and GET jointly model RNA and ATAC modalities for regulatory inference. Grounding leverages atlas-derived cCRE annotations and transcription-factor motifs. Performance is strong for biological grounding and multi-omics alignment, but interpretability is limited to motif attention visualization.
C.5 Agentic Frameworks (Extended Details)
Agentic models couple (i) a pretrained molecular encoder, (ii) an LLM-based controller, and (iii) external tool interfaces. scAgent executes multi-step annotation workflows via ontology queries; CellVerse coordinates multiple domain agents for transcriptomic, spatial, and literature reasoning; EpiAgent extends agentic control to enhancer–gene analysis; and Teddy/scPilot (Joodaki et al., 2024) provide lightweight orchestrators for benchmarking pipelines. Grounding sources include the Human Cell Atlas ontology, UBERON, and PubMed. These systems score highest in Explainability and Emerging paradigms, marking a transition from static embeddings to interactive, goal-directed modeling. Open challenges include evaluation of reasoning accuracy, privacy, and reproducibility.
Appendix D Datasets
Table 1: RNAseq single-cell datasets used in LLM-based single-cell research.
Dataset Tasks Description Scale Link / Citation
Tabula Sapiens v2 Annotation, Integration Multi-organ human atlas ($\sim$1.1 M cells, 28 organs across 24 donors) capturing cell-type heterogeneity and cross-tissue transcriptional variation; droplet and plate-seq modalities. 1.1 M cells / 28 organs / 24 donors CZ Biohub Portal
Tabula Muris (mouse, multi-organ) Annotation, Integration Mouse multi-organ atlas of gene expression combining droplet and FACS; enables cross-tissue comparison and batch integration. $\sim$100,000 cells / 20 organs HCA Project Page
Mouse Cell Atlas (scMCA) Annotation, Integration Comprehensive mouse single-cell atlas constructed with Microwell-seq; supports cell-type matching via scMCA tool. >>400,000 cells / >>40 tissues MCA Portal (ZJU)
Human Lung Cell Atlas (HLCA v1.0) Annotation, Integration, Disease Modeling Integrated human lung reference built from 49 scRNA-seq datasets ($\sim$2.4 M cells) across 16 studies; unified epithelial, immune, and stromal annotations. $\sim$2.4 M / 49 datasets / 16 studies HLCA v1.0 Portal
HCA lung project (example project page) Annotation, Integration, Disease Modeling HCA lung cohort integrating $\sim$2.4 M single cells from 49 datasets; harmonized annotations for airway, immune, and endothelial populations. 2.4 M cells / 49 datasets HLCA Project Page
Allen Brain Map – Cell Types RNA-seq (human & mouse) Annotation, Cross-species, Trajectory Single-cell and single-nucleus transcriptomes from human and mouse brain regions; supports cross-species comparison and cell-type taxonomy benchmarking. >>1.8 M cells / multiple cortical & subcortical regions Allen Brain Cell Types Database
Yale Lung Disease Cell Atlases (e.g., COPD) Annotation, Disease Modeling Single-cell RNA atlases of diseased human lungs (COPD, IPF) from Yale’s Kaminski lab, capturing altered epithelial, endothelial, and immune populations. $\sim$300k cells (IPF + COPD + controls) HCA/Yale Atlas
Table 2: ATACseq and chromatin-accessibility datasets used in LLM-based single-cell research.
Dataset Tasks Description Scale Link / Citation
Mouse sci-ATAC-seq atlas (Cusanovich et al., 2018) Annotation, Trajectory, GRN inference Landmark single-cell ATAC-seq atlas profiling $\sim$100,000 nuclei across 13 adult mouse tissues using combinatorial indexing (sci-ATAC-seq); enables cross-tissue regulatory and lineage analysis. $\sim$100k nuclei / 13 tissues GSE111586, Science 2018
Human adult scATAC atlas (Zhang et al., 2021) Annotation, GRN, Cross-tissue integration Comprehensive single-cell chromatin accessibility atlas of adult human tissues, defining candidate cis-regulatory elements (cCREs) across 25 tissues and 222 cell types; foundation of the ENCODE human cCRE registry. $\sim$472k nuclei / 25 tissues Nature 2021, ENCODE Portal
Human fetal scATAC atlas (Domcke et al., 2020; GSE149683) Annotation, Developmental trajectory, GRN Single-cell ATAC-seq atlas of human fetal tissues profiling 15 organs across mid-gestation; reveals developmental enhancer activity and regulatory lineage trajectories. $\sim$720k nuclei / 15 organs Science 2020, GSE149683
Massively parallel scATAC (Satpathy et al., 2019) Annotation, Trajectory, Regulatory modeling Pioneering high-throughput scATAC-seq dataset of $\sim$200k immune and cancer nuclei enabling scalable mapping of open-chromatin landscapes; forms benchmark for lineage and immune-cell trajectory studies. $\sim$200k nuclei / blood and tumor tissues Nat. Biotechnol. 2019, GSE123581
ENCODE portal (single-cell experiments) Annotation, Integration, Regulatory modeling Centralized repository from the ENCODE Consortium aggregating thousands of single-cell RNA and ATAC assays from human and mouse tissues; provides uniformly processed metadata, peak calls, and cCRE annotations for benchmarking. >>2,000 single-cell assays / multiple tissues ENCODE Portal
T cell epigenetic atlas (Giles et al., 2022) Trajectory, GRN, Disease modeling Single-cell ATAC-seq atlas profiling chromatin accessibility across human T-cell activation, exhaustion, and differentiation states; supports trajectory reconstruction and immune-epigenetic modeling. $\sim$150k nuclei / T-cell subsets Nat. Immunol. 2022
Table 3: Multiome datasets (paired/tri-omic: RNA + ADT, RNA + ATAC, ATAC + ADT, and tri-modal) used in LLM-based single-cell research.
Dataset Tasks Description Scale Link / Citation
UCSC Cell Browser Hub Annotation, Integration, Visualization Aggregated repository of hundreds of public single-cell datasets across species and modalities with metadata and embeddings; useful for exploratory analysis and reference selection. >>1,200 datasets / multi-species UCSC Cell Browser Hub
Human Cell Atlas (HCA) data browser (multi-project) Annotation, Integration, Cross-project mapping A global human cell atlas aggregating $\sim$63.2 M cells across 515 projects and >>11,000 donors, spanning diverse tissues and modalities; enables cross-study reference and meta-integration. $\sim$63.2 M cells / 515 projects / 11,000++ donors HCA Data Portal
Azimuth reference collections (PBMC, lung, kidney, fetal) (RNA + ADT) Annotation, Integration Curated single-cell reference atlases by the Satija Lab for automated cell-type mapping in Seurat/Azimuth; harmonized labels and multimodal ADT features. 100k–1 M cells across multiple organs Azimuth Data Portal
TEA-seq (tri-modal RNA + ATAC +  ADT; GSE158013) Integration, Perturbation, Multi-modal reasoning Tri-modal PBMC dataset measuring transcriptome, chromatin accessibility, and surface proteins; supports alignment/fusion of omics layers and pretraining. $\sim$100k cells GSE158013
DOGMA-seq (RNA + ATAC +  Protein) Integration, Perturbation, Cross-modal reasoning Tri-modal profiling of transcriptome, chromatin, and proteins in human immune cells; benchmark for joint embeddings and cross-modal translation. $\sim$50k cells / PBMCs Nature Cell Biol. 2022 / GSE184715
ASAP-seq (ATAC + Protein) Integration, GRN, Epigenetic modeling Paired chromatin accessibility and surface proteins via antibody-derived tags; links cis-regulatory variation with immune phenotypes. $\sim$100k nuclei / PBMCs Nat. Biotechnol. 2021 / GSE162690
CITE-seq compendia (RNA + ADT) Annotation, Integration, Transfer learning Large compendium of paired RNA and surface-protein profiles across blood and tissue; enables multimodal foundation pretraining and zero-shot annotation. >>500k cells across multiple tissues Nat. Methods 2019 / GSE100866
Multiome Benchmark Pack (QuKun Lab) Integration, Scalability, Batch-effect analysis Public benchmark suite consolidating 25 RNA + Protein, 12 RNA + ATAC, and 4 tri-omic datasets (CITE/TEA/DOGMA), standardized for cross-modal and large-scale integration testing. 41 datasets total Benchmark Portal
Table 4: Spatial transcriptomics and imaging datasets used in LLM-based single-cell research.
Dataset Tasks Description Scale Link / Citation
Slide-seq / Slide-seqV2 Spatial mapping, Trajectory, Integration High-resolution spatial transcriptomics of mouse brain (hippocampus, cerebellum, cortex) at $\sim$10 µm bead resolution; supports neighborhood and spatial-domain reconstruction benchmarks. $\sim$50k–100k spots per tissue Nat. Biotechnol. 2020 / SCP815
MERFISH (imaging) Spatial mapping, Pathway, Annotation Multiplexed error-robust fluorescence in situ hybridization (MERFISH) datasets profiling millions of mouse-brain cells with 3D coordinates and 483–1,000-gene panels; benchmark for high-resolution spatial mapping. $\sim$4 M cells / 483–1,000 genes Vizgen MERFISH Portal / Science 2018
Stereo-seq (STOmics) Spatial mapping, Developmental trajectory Genome-scale Stereo-seq datasets with submicron resolution; includes MOSTA (mouse embryo) and 3D Drosophila atlases for developmental and cross-species modeling. $\sim$100 µm $\times$ 100 µm tiles / millions of spots Cell 2022 / STOmics Data Hub
DBiT-seq & spatial multi-omics Spatial mapping, Integration, Multi-omic reasoning Deterministic barcoding in tissue (DBiT-seq) capturing spatial RNA and protein expression in mouse embryo and human lymph node; supports multi-omic spatial benchmarks. $\sim$20k spots / 100 µm grids Nat. Biotechnol. 2020 / GSE152506
10x Visium / Visium HD Spatial mapping, Annotation, Integration Widely used capture-array platform for spatial transcriptomics; Visium uses 55 µm spots, Visium HD extends to 2 µm grids with improved gene recovery and FFPE compatibility. $\sim$5k–55k spots per tissue 10x Genomics Visium Portal
Xenium / CosMx (in situ) Spatial mapping, Clinical translation, Privacy In situ spatial transcriptomics platforms (10x Xenium, NanoString CosMx) profiling thousands of transcripts in human FFPE and fresh-frozen tissues (e.g., breast, colon, NSCLC); bridge omics and histology for clinical translation. 10k–100k cells per section Xenium Explorer  /  CosMx Portal
Table 5: Perturbation and Drug-Response single-cell datasets used in LLM-based frameworks.
Dataset Tasks Description Scale Link / Citation
Genome-scale Perturb-seq (Replogle 2022) processed data Perturbation, GRN, Causal inference Largest CRISPR-based Perturb-seq dataset profiling $\sim$2.5 M single cells across >>2 000 genetic perturbations; benchmark for causal network inference and representation learning. $\sim$2.5 M cells / >>2 000 perturbations Cell 2022 / Figshare Dataset
Norman et al. 2019 Perturb-seq Perturbation, Combinatorial GRN Foundational combinatorial CRISPR Perturb-seq dataset (immune + cancer models); establishes feasibility of pooled functional genomics at single-cell resolution. $\sim$200 k cells / >>250 combinations Science 2019 / GSE133344
Dixit et al. 2016 (GSE90063) Perturbation, Combinatorial GRN Early Perturb-seq study targeting 24 genes in macrophages; establishes pipeline for pooled CRISPR screens with single-cell RNA-seq. $\sim$100 k cells / 24 target genes Cell 2016 / GSE90063
Adamson et al. 2016 Perturbation, Stress-response, GRN Single-cell CRISPR Perturb-seq of ER-stress pathways in K562 cells; benchmark for pathway-level perturbation responses. $\sim$30 k cells / 10 perturbations Cell 2016 / GSE90060
sci-Plex collection (drug screens) + cellxgene portal Perturbation, Stress-response, GRN Multiplexed chemical-perturbation scRNA-seq of >>650 k cells treated with 188 compounds via sci-Plex barcoding; pharmacotranscriptomic profiles for drug response modeling. $\sim$650 k cells / 188 drugs Cell 2020 / cellxgene Collection
Compressed Perturb-seq (immune LPS, 598 genes) Drug response, Perturbation, Benchmarking Low-multiplicity Perturb-seq library targeting 598 immune genes in macrophages under LPS stimulation; enables compressed experimental design for causal modeling. $\sim$300 k cells / 598 targets bioRxiv 2021 / GSE179924
In vivo Perturb-seq (brain / ASD genes) Perturbation, GRN, Neuroscience In vivo CRISPR Perturb-seq targeting $\sim$30 ASD-linked genes in mouse cortex; maps neuronal gene-regulatory networks in native contexts. $\sim$200 k cells / 30 genes Science 2023 / GSE216113
Virtual Cell Challenge PBMC cytokine perturbations Perturbation, GRN, Benchmarking Community benchmark dataset (ARC Institute 2024) of $\sim$300 k PBMC cells under cytokine stimulation and CRISPRi perturbations; standardized splits for LLM and state-transition evaluation. $\sim$300 k cells / 150 perturbed genes ARC Challenge Repository / Kaggle
Table 6: Plant RNA single-cell transcriptomics datasets used in LLM research.
Dataset Tasks Description Scale Link / Citation
scPlantDB (meta-collection) Annotation, Integration, Meta-analysis Comprehensive plant single-cell transcriptome database integrating 67 datasets from 17 plant species ($\sim$2.5 M cells); unified preprocessing and annotations; supports cross-species modeling and plant-specific LLMs. $\sim$2.5 M cells / 67 datasets / 17 species scPlantDB Portal
PlantscRNAdb Annotation, Marker discovery Curated database of plant cell-type marker genes across four species (Arabidopsis, rice, maize, tomato); supports ontology-based annotation and cell identity benchmarking. 4 species / multiple tissues PlantscRNAdb Portal
Arabidopsis scRNA-seq (E-CURD-4) Annotation, Trajectory, Development Baseline scRNA-seq dataset of Arabidopsis thaliana root and leaf tissues ($\sim$10,779 cells); used for developmental lineage and differentiation analysis. 10,779 cells / 2 tissues EBI ArrayExpress E-CURD-4
Tobacco leaf scRNA-seq (transgenic antibody line) Annotation, Perturbation, Stress response Single-cell transcriptomic profiling of transgenic Nicotiana tabacum leaves expressing llama antibody; identifies immune-like responses and cell-type heterogeneity under genetic perturbation. $\sim$25k cells / leaf tissue Sci. Data 2023
Plant scRNA Browser (PscB) Annotation, Visualization, Cross-tissue integration Online visualization hub aggregating plant scRNA-seq datasets (Arabidopsis, rice, Wolffia) with harmonized metadata and interactive UMAP-based cell-type search. 15+ datasets / 3+ species PscB Data Portal
Appendix E Methods Comparison
Table 7: Full comparison of single-cell LLM and agentic methods (Domains column removed). Year merged into Model.
Model (Year) Published where Category Modality Grounding Type Agentic (Y/N) Primary Task Domain Score
scGPT Cui et al. (2024) Nature Methods Foundation Multiomics (scRNA + optionally multi-omic modes) Atlas (trained on large atlas of single-cell datasets) No Annotation, Integration, Perturbation (annotation as primary) 6
Geneformer Theodoris et al. (2023) Nature Foundation scRNA None / Rank-based (implicit) No Cell classification, in-silico perturbation, network prediction 6
scFoundation Hao et al. (2024) Nature Methods Foundation scRNA / multi-omics (transcriptomics focus) Atlas (large pretrained corpus) No Cell annotation, perturbation prediction, drug response, gene module inference 7
CellFM Zeng et al. (2025) Nature Communications Foundation scRNA Atlas / value-projection No Learns embeddings from 100M cells and supports annotation, perturbation, gene function 7
iSEEEK Shen et al. (2022) Briefings in Bioinformatics Foundation scRNA Atlas (11.9M cells, human + mouse) No Integrates massive single-cell datasets via gene-ranking similarity to enable scalable cross-dataset embedding and clustering 5
tGPT Shen et al. (2023) iScience Foundation Model (Autoregressive) scRNA Atlas (22.3 M cells) No Treats gene-expression ranks as token sequences and learns generative embeddings for cell clustering, trajectory inference, and bulk-tissue analysis 5
scBERT Yang et al. (2022) Nature Machine Intelligence Foundation Model (Encoder-only) scRNA Atlas (public scRNA-seq corpora) No Learns bidirectional gene–cell embeddings using BERT-style masked modeling for robust cell-type annotation and novel cell discovery 4
UCE (Universal Cell Embeddings) Rosen et al. (2023) BioRxiv Foundation scRNA (cross-species) Atlas / binary masked self-supervision No Embeds any cell zero-shot across species into a shared latent space for clustering, lineage inference, annotation 6
GeneCompass Yang et al. (2024b) Nature-/Cell Research Foundation Model scRNA / cross-species Ontology (GRN + co-expression + gene family knowledge) No Learns cross-species gene–cell embeddings and supports annotation, perturbation, dose-response, and GRN inference 6
scELMo Liu et al. (2023) BioRxiv Text-Bridge LLM scRNA + Metadata Text LLM-generated embeddings from gene/metadata descriptions + raw data No Embeds cells via text-derived gene metadata embeddings combined with expression, then supports clustering, batch correction, annotation, perturbation 5
CellPLM Wen et al. (2023) ICLR Spatial (Multimodal Foundation Model) scRNA + Spatial (SRT) Atlas + Spatial relations (uses SRT in pretraining) No Treats cells as tokens and tissues as sentences, leveraging spatially-resolved transcriptomics and a Gaussian-mixture prior to encode inter-cell relations for denoising, spatial imputation, and perturbation prediction. 4
scMoFormer Tang et al. (2023) ArXiv Spatial (Multimodal Foundation Model) scRNA + Protein / Multi-omics Atlas + domain knowledge in cross-modality aggregation No Uses modality-specific transformers and cross-attention to impute missing modalities, classify cells, and fuse multimodal representations 4
scFormer Xu et al. (2024) ArXiv Spatial (Multimodal Foundation Model) (or Foundation + multi-omics) Transcript- omics + Proteomics / multimodal Atlas + transformer-based fusion No Aligns and integrates multi-omics single-cell data, recovers missing modalities, and transfers labels across modalities 4
scMulan Bian et al. (2024a) BioRxiv Foundation Model scRNA + Metadata Atlas + prompt-conditioned generative modeling No Encodes each cell as a “c-sentence” integrating expression + metadata; supports zero-shot annotation, batch integration, and conditional generation 3
scPRINT Kalfon et al. (2025) Nature Communications Foundation Model scRNA Atlas (50M+ cell pretraining) No Learns cell embeddings and infers cell-specific gene regulatory networks; supports zero-shot denoising, batch correction, label prediction, expression reconstruction 5
scGraphformer Fan et al. (2024) Nature Communications Biology Foundation Model scRNA Atlas + relational prior (kNN bias, refined) No Learns a cell–cell graph via a transformer-GNN hybrid for better classification and interaction inference 4
scRDiT Dong et al. (2025) ArXiv Foundation Model scRNA (transcriptome) Atlas-like generative prior / diffusion modeling No Generates synthetic scRNA-seq samples via diffusion transformer + DDIM for accelerated sampling 2
scGFT Nouri (2025) Nature Communications Biology Foundation Model scRNA Fourier-based perturbation / reconstruction (train-free) No Synthesizes new single-cell expression profiles by perturbing Fourier components in frequency space. 4
scTrans Lu et al. (2024) IJCAI Foundation Model scRNA Sub-vector masked completion over gene modules No Learns multi-scale sub-vector tokens to perform gene-selective cell-type annotation via masked completion and contrastive regularization 3
scGT (Graph Transformer) Qi et al. (2025b) Bioinfo-rmatics Advance Graph / Multi-omics Integration scRNA + scATAC Observed / Hybrid Graph No Multi-omics integration + label transfer 4
TransformerST Lu et al. (2024) Briefings in Bioinformatics Spatial / Multimodal FM Spatial (histology + gene expression) Spatial (histology + gene expression) No Super-resolution gene expression & tissue clustering 3
scGPT-spatial Cui et al. (2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial Transcriptomics + scRNA prior Atlas + spatial-aware decoding No Extends scGPT via continual pretraining to spatial data, supports multi-slide integration, cell-type deconvolution, and spatial gene imputation 3
HEIST Madhu et al. (2025) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + proteomics Hierarchical graph modeling (spatial + GRNs) No Learns joint embeddings of cells and genes in spatial context to perform cell annotation, gene imputation, spatial clustering, clinical outcome prediction 4
stFormer Cao et al. (2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + ligand context Atlas + biased cross-attention to ligand niche genes No Learns gene embeddings contextualized by ligand signals in spatial microenvironments; aids clustering, ligand–receptor inference, and perturbation simulation 4
FmH2ST Wang et al. (2025) Nucleic Acids Research Spatial (Multimodal Foundation Model) Histology image + spatial transcriptomics Image foundation + dual graphs + spot branch No Predict spatial gene expression from histology using fused image and spot features, supporting denoising, heterogeneity detection, and regulatory inference 3
OmiCLIP Cui et al. (2025) Nature Methods Spatial (Multimodal Foundation Model) Histology + Spatial Transcriptomics Image–gene contrastive alignment (rank-based) No Learns unified visual-omics embeddings linking histopathology and spatial gene expression; enables cross-modal prediction, annotation, and tissue retrieval 7
QuST-LLM Huang (2024) ArXiv Text-Bridge LLM Spatial transcriptomics + histology metadata GO-term + gene enrichment + LLM narrative overlay Yes Converts ST data and ROIs into human-readable narratives and matches natural language queries to spatial regions via GO/LLM interpretation 4
SpaCCC Yang et al. (2024a) IEEE Xplore Text-Bridge LLM Spatial transcriptomics / transcriptomic genes (LRs) LLM embeddings of ligand + receptor genes No Infers spatially resolved cell–cell communication by embedding LR pairs in LLM latent space + diffusion / permutation test filtering 3
spaLLM Ji et al. (2024) Briefings in Bioinformatics Spatial (Multimodal Foundation Model) Spatial multi-omics (RNA, ATAC, protein) scGPT embeddings + GNN + attention fusion No Enhances spatial domain identification by fusing LLM-derived embeddings and spatial-omics signals via multi-view attention 3
scMMGPT (2025) Shi et al. (2025) ArXiv Text-Bridge LLM / Multimodal_FM RNA + Text (metadata) Annotation, description No Generation - links single-cell and text PLMs to describe cells, generate pseudo-cells from text, and enhance annotation through text-conditioned reasoning 6
Cell2Sentence (C2S) Levine et al. (2024) PMLR Text-Bridge LLM RNA (scRNA-seq converted to “cell sentences”) Marker / Implicit (gene rank-order) No Generation, annotation & reconstruction - encodes cells as gene-ranked sentences and fine-tunes LLMs to classify or generate biologically meaningful cell text 6
C2S-Scale Rizvi et al. (2025) BioRxiv Text-Bridge LLM+ Multimodal Foundation Model RNA + Text / Metadata Atlas + Text / Implicit rank grounding No Chat, generation & annotation - trains large LLMs on cell sentences and biological text to enable perturbation prediction and multicellular summarization 7
Cell2Text Kharouiche et al. (2025) ArXiv Text-Bridge LLM + Multimodal Foundation Model RNA → Natural Language Gene-level embeddings + ontology / metadata grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict expression and interpret TF–cis interactions 7
GenePT Chen and Zou (2024) BioRxiv Foundation Model / Text-Augmented scRNA Literature / Text embedding grounding No Embedding & downstream prediction - uses GPT-3.5 gene text embeddings to derive cell embeddings via weighted or ranked gene aggregation. 5-7
CellLM Zhao et al. (2023) ArXiv Foundation Model / Representation model RNA (single cell expression) Implicit - contrastive embedding from expression data No Represent cells via a contrastive-learning transformer, optimize embedding space for tasks like cell-type annotation, drug sensitivity prediction 6
scExtract Wu and Tang (2025) Genome Biology Agentic Framework (Text-Bridge LLM) RNA (scRNA-seq) + Text Article-based parameter extraction Yes Annotation & Integration - uses LLMs to extract pipeline parameters from publications and apply them for dataset harmonization. 7
scAgent Mao et al. (2025) ArXiv Agentic Framework (Text-Bridge hybrid) RNA (scRNA-seq) Reference atlas + marker gene reasoning + memory grounding Yes Universal cell annotation & novel cell discovery - scAgent uses an LLM planning module, memory, and tool modules to annotate cells across tissues, detect unknown types, and incrementally learn new annotations 7
EpiFoundation Wu et al. (2025) BioRxiv Epigenomic Foundation Model / Multimodal alignment scATAC (chromatin accessibility) Peak-to-gene supervision (alignment to expression) No Cell embedding, annotation, batch correction, gene expression prediction - trains on sparse peak sets, aligns to gene expression supervision, and transfers learned embeddings for downstream ATAC tasks 7
EpiAgent Chen et al. (2025b) BioRxiv Epigenomic Foundation Model / Multimodal text-bridge hybrid scATAC / chromatin accessibility Peak tokenization + external embeddings + regulatory supervision grounding No Embedding, annotation, imputation, and perturbation prediction - encodes chromatin accessibility as ranked cCRE tokens, enabling zero-shot cell annotation, peak imputation, and response prediction. 8
ChromFound Jiao et al. (2025) Nature Epigenomic → Expression Transformer / Regulatory FM scATAC (chromatin accessibility) + DNA sequence Motif × peak matrix / masked regulatory grammar grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict gene expression in seen and unseen cell types; also interprets transcription factor interactions and cis-regulatory elements 8
GET (General Expression Transformer) Fu et al. (2025) Nature Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
scMPT Palayew et al. (2025) ArXiv Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
EpiBERT Javed et al. (2025) Cell Genomics Epigenomic Foundation Model / Multi-modal Transformer DNA sequence + chromatin accessibility Masked-accessibility pretraining + motif & sequence fusion No Accessibility imputation, gene expression prediction & regulatory inference - predicts masked ATAC signals, then fine-tunes to predict expression and enhancer-gene links, generalizing to unseen cell types 8
scMamba Yuan et al. (2025) ArXiv Multimodal / Foundation_FM Multi-omics (RNA + others) Implicit via integrated features; no explicit external grounding No Multi-omic embedding, integration & annotation across modalities without prior feature selection 7
GeneMamba Qi et al. (2025a) ArXiv Foundation Model / State-Space Model RNA (scRNA) Implicit (via gene sequence context + pathway loss) No Multi-batch integration, cell-type annotation, gene correlation - uses a BiMamba state-space architecture with pathway-aware losses for scalable, context-rich modeling of scRNA 7
Nicheformer Schaar et al. (2024) BioRxiv Foundation _FM/ Spatial / Multimodal_FM scRNA + spatial transcriptomics Contextual tokenization + metadata + spatial neighborhood embedding No Spatial context prediction, spatial label / niche prediction, mapping spatial info to dissociated cells 7
scFormer Cell+ Cui et al. (2022) Bioarxiv Foundation Model / Multimodal Transformer RNA (+ metadata) Joint gene-cell embedding with metadata tokens No Integration & Annotation - context-aware joint embedding for cross-species/tissue generalization 7
scPlantLLM Cao et al. (2025) Genomics, Proteomics and Bioinformatics Text-Bridge LLM (Foundation Model) RNA (plant scRNA-seq) Gene token + binned expression embedding (Plant Atlas grounding) No Annotation, Integration, GRN inference - pretrains on plant single-cell data with masked LM and cell-type supervision, enabling cross-species annotation, clustering, and regulatory discovery in plant systems 7
LICT Ye et al. (2024) ArXiv Text-Bridge LLM / Annotation LLM Hybrid RNA (scRNA-seq) Marker gene–based DE lists + LLM prompting No Annotation & Reliability - iterative LLM prompting with DE markers for label refinement and confidence scoring 5
Teddy (family of models) Chevalier et al. (2025) ArXiv Foundation Model / Disease-aware Transformer scRNA (single-cell RNA-seq) Self-supervised + supervised annotation supervision No Disease state classification / healthy vs diseased detection - trained to identify disease conditions of held-out donors and distinguish diseased vs healthy cells in new disease contexts 7
Pilot Joodaki et al. (2024) Github Benchmark / Evaluation Framework RNA Model-agnostic pilot foundation testing No Benchmarking & Evaluation - lightweight pilot framework for early single-cell FM testing 4
CellVerse Zhang et al. (2025b) ArXiv Benchmark / Evaluation Multi-omics (RNA, CITE, ASAP, etc.) Implicit via prompt encoding No QA benchmark for annotation, drug response, perturbation tasks 5
xTrimoGene Gong et al. (2023) NeurIPS / ArXiv Foundation Model (Scalable Transformer) scRNA-seq Sparse masking + auto-discretization No Representation learning + annotation, perturbation, drug synergy prediction 7
EpiAttend Li et al. (2022) NeurIPS 2022 Workshop Regulatory / Sequence-Epigenome Transformer DNA sequence + single-cell epigenomic data Sequence + cell-specific epigenome grounding No Predict cell type–specific gene expression by integrating DNA sequence and single-cell epigenomic tracks, linking enhancers and promoters 6
Spatial2-Sentence Chen et al. (2025a) ArXiv Spatial / Text-Bridge hybrid Spatial + expression (Imaging Mass Cytometry) Spatial adjacency + expression similarity tokenization No Encode spatial & expression context into multi-sentence prompts for LLMs to perform cell-type classification and clinical status prediction 6
ChatCell Fang et al. (2024) ArXiv Text-Bridge LLM (Instructional) scRNA → “cell sentence” Vocabulary adaptation + unified sequence generation of cell sentences No Natural language interface for single-cell tasks - allows users to query, annotate, generate, and explore scRNA data via text prompts. Hugging Face 6
CellAtria Nouri et al. (2025) BioRxiv Agentic Framework scRNA- seq + metadata Ontology (graph + metadata) Yes Annotation & Ontology Mapping 7
CellAgent Xiao et al. (2024) ArXiv Agentic Framework scRNA- seq None Yes Annotation & Ontology Mapping 7
Appendix F Appendix Figure
Figure 3: Task vs Model Heatmap
Figure 4: Comparison of task coverage between agentic (dark blue) and non-agentic (yellow) models. Agentic frameworks emphasize annotation, ontology mapping, spatial mapping while non-agentic models concentrate on trajectory, perturbation modeling, regulatory and pathway inference as well.
Figure 5: Comparison of domain coverage between agentic (dark blue) and non-agentic (yellow) models. Agentic frameworks emphasize explainability, fairness, and emerging paradigms, while non-agentic models concentrate on biological grounding and batch effects.

### v2

@@
 Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text.
-In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework.
+(Zhang et al., 2025a; Aljohani et al., 2025; Bi et al., 2024).
+Prior work on computational workflows and biological prediction tasks demonstrates how learned representations capture complex molecular relationships and enhance practical performance.
+(Lam et al., 2024; Bhattacharya et al., 2024; Dip et al., 2024; Nam et al., 2024; Shuvo et al., 2024) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework.
 Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis.
@@
 Despite this progress, the field remains fragmented.
-Models differ widely in data modality, supervision type, and evaluation standard—ranging from scRNA-seq to ATAC, multi-omic, and spatial measurements—hindering cross-model comparison and reproducibility.
+Models differ widely in data modality, supervision type, and evaluation standard ranging from scRNA-seq to ATAC, multi-omic, and spatial measurements hindering cross-model comparison and reproducibility.
 Benchmarking is inconsistent, datasets are unevenly distributed across modalities, and agentic frameworks lack standardized ways to measure reasoning correctness or biological grounding.
@@
 To address these challenges, we introduce LLM4Cell, a comprehensive survey of large language and agentic models for single-cell biology.
-We analyze 58 representative methods spanning five methodological families—Foundation, Text-Bridge, Spatial/Multimodal, Epigenomic, and Agentic—and organize them across eight major tasks, from annotation and trajectory inference to perturbation and drug-response modeling.
+We analyze 58 representative methods spanning five methodological familiesFoundation, Text-Bridge, Spatial/Multimodal, Epigenomic, and Agentic and organize them across eight major tasks, from annotation and trajectory inference to perturbation and drug-response modeling.
 Complementing this, we compile over 40 publicly available datasets across RNA, ATAC, multi-omic, spatial, perturbation, and plant domains, and evaluate each model using a ten-dimension rubric covering grounding, fairness, scalability, and interpretability.
 Contributions.
-LLM4Cell provides: (1) a modality-balanced registry of $\sim$40 benchmark datasets for large-model training and evaluation; (2) a unified taxonomy of 58 foundation and agentic models across eight analytical tasks; (3) a ten-dimension domain rubric capturing biological grounding, generalization, and ethics; and (4) a critical discussion of open challenges in cross-modal alignment, reasoning, and trustworthy AI for cell biology.
+LLM4Cell provides: (1) a modality-balanced registry of around 40 benchmark datasets for large-model training and evaluation; (2) a unified taxonomy of 58 foundation and agentic models across eight analytical tasks; (3) a ten-dimension domain rubric capturing biological grounding, generalization, and ethics; and (4) a critical discussion of open challenges in cross-modal alignment, reasoning, and trustworthy AI for cell biology.
 2 Related Work
@@
 However, existing reviews remain limited in scope.
-They focus on architectures or prompting but give little attention to the datasets underpinning model training and evaluation—especially those covering spatial, perturbation, or multimodal modalities.
-Standardized benchmarks, domain-aware metrics, and reproducible splits are seldom addressed, and plant single-cell datasets—key for cross-kingdom generalization—are almost entirely absent.
+They focus on architectures or prompting but give little attention to the datasets underpinning model training and evaluation especially those covering spatial, perturbation, or multimodal modalities.
+Standardized benchmarks, domain-aware metrics, and reproducible splits are seldom addressed, and plant single-cell datasets key for cross-kingdom generalization are almost entirely absent.
 Moreover, agentic reasoning and tool-augmented systems are typically treated as isolated examples rather than a coherent methodological class, leaving the relationship among data modalities, modeling paradigms, and biological domains fragmented.
@@
 Progress in large language models for single-cell biology relies on the rapid growth of high-quality datasets across transcriptomic, epigenomic, multimodal, and spatial domains.
-Our survey compiles over forty public resources spanning five major modalities—RNA, ATAC, multi-omic, spatial, and perturbation/drug-response—plus emerging plant single-cell atlases.
+Our survey compiles over forty public resources spanning five major modalities RNA, ATAC, multi-omic, spatial, and perturbation/drug-response plus emerging plant single-cell atlases.
 These datasets underpin model pretraining, evaluation, and cross-modal reasoning for annotation, integration, trajectory inference, perturbation modeling, and spatial mapping.
@@
 Large language models are rapidly reshaping single-cell biology, producing diverse architectures, pretraining schemes, and reasoning frameworks.
-We categorize 58 representative methods into five methodological families—Foundation Models, Text-Bridge LLMs, Spatial and Multimodal Models, Epigenomic Models, and Agentic Frameworks—based on their core design and data modality.
+We categorize 58 representative methods into five methodological families Foundation Models, Text-Bridge LLMs, Spatial and Multimodal Models, Epigenomic Models, and Agentic Frameworks based on their core design and data modality.
 These span the progression from gene-level embeddings to multimodal and autonomous systems capable of biological reasoning.

### v3

@@
 Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text.
-(Zhang et al., 2025a; Aljohani et al., 2025; Bi et al., 2024).
+(zhang2025survey; aljohani2025comprehensive; bi2024ai).
 Prior work on computational workflows and biological prediction tasks demonstrates how learned representations capture complex molecular relationships and enhance practical performance.
-(Lam et al., 2024; Bhattacharya et al., 2024; Dip et al., 2024; Nam et al., 2024; Shuvo et al., 2024) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework.
-Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis.
+(lam2024large; bhattacharya2024large; dip2024deepage; nam2024using; shuvo2024assessing) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework.
+Early foundation models such as scGPT  (cui2024scgpt), Geneformer  (theodoris2023transfer), and scFoundation  (hao2024large) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (zhao2023large), scAgent  (mao2025scagent), and CellVerse  (zhang2025cellverse) extend these capabilities toward reasoning, dialogue, and autonomous analysis.
 Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
@@
 The intersection of large language models (LLMs) and single-cell biology has gained rapid momentum, motivating several early surveys and benchmarks.
-The ACL 2025 survey on foundation models for single-cell biology (Zhang et al., 2025a) outlines pretrained and fine-tuned architectures across protein, gene, and cell representations.
-Broader reviews such as Lan et al.
-(2025) and Bian et al.
-(2024b) discuss “large cellular models” (LCMs) and trends in scaling, tokenization, and transfer learning across biological data.
-Benchmark studies including Single-Cell Omics Arena (Liu et al., 2024) and CellVerse (Zhang et al., 2025b) evaluate LLMs on annotation and question answering, while works like scInterpreter (Guo et al., 2023) and scExtract (Wu and Tang, 2025) demonstrate practical integration into annotation and curation pipelines.
+The ACL 2025 survey on foundation models for single-cell biology (zhang2025survey) outlines pretrained and fine-tuned architectures across protein, gene, and cell representations.
+Broader reviews such as lan2025large and bian2024general discuss “large cellular models” (LCMs) and trends in scaling, tokenization, and transfer learning across biological data.
+Benchmark studies including Single-Cell Omics Arena (liu2024single) and CellVerse (zhang2025cellverse) evaluate LLMs on annotation and question answering, while works like scInterpreter (guo2023scinterpreter) and scExtract (wu2025scextract) demonstrate practical integration into annotation and curation pipelines.
 However, existing reviews remain limited in scope.
@@
 These datasets underpin model pretraining, evaluation, and cross-modal reasoning for annotation, integration, trajectory inference, perturbation modeling, and spatial mapping.
-Transcriptomic atlases such as Tabula Sapiens, Tabula Muris  (Consortium* et al., 2022), and the Human Cell Atlas  (Travaglini et al., 2020) remain dominant for foundation-model training and ontology-aware annotation.
+Transcriptomic atlases such as Tabula Sapiens, Tabula Muris  (tabula_sapiens2022), and the Human Cell Atlas  (travaglini2020_hlca) remain dominant for foundation-model training and ontology-aware annotation.
 Chromatin-accessibility data (e.g.
-Cusanovich mouse atlas (Cusanovich et al., 2018), human adult/fetal scATAC (Domcke et al., 2020)) map regulatory states but remain sparse and heterogeneous.
+Cusanovich mouse atlas (cusanovich2018_science), human adult/fetal scATAC (zhang2021_encode)) map regulatory states but remain sparse and heterogeneous.
 Multi-omic resources (e.g.
-TEA-seq  (Swanson et al., 2021), DOGMA-seq  (Mimitou et al., 2021), CITE-seq (Stoeckius et al., 2017)) link RNA, ATAC, and protein modalities, providing supervision for cross-view alignment.
+TEA-seq  (swanson2021_natbiotech), DOGMA-seq  (mimitou2021scalable), CITE-seq (stoeckius2017simultaneous)) link RNA, ATAC, and protein modalities, providing supervision for cross-view alignment.
 Spatial technologies (e.g.
-Visium  (Oliveira et al., 2025), Slide-seqV2  (Stickels et al., 2021), MERFISH  (Chen et al., 2015), Stereo-seq  (Chen et al., 2022)) connect molecular profiles to tissue architecture, while functional datasets (e.g.
-Perturb-seq (Replogle et al., 2022)) benchmark causal reasoning.
+Visium  (oliveira2025high), Slide-seqV2  (stickels2021highly), MERFISH  (chen2015spatially), Stereo-seq  (chen2022spatiotemporal)) connect molecular profiles to tissue architecture, while functional datasets (e.g.
+Perturb-seq (replogle2022_cell)) benchmark causal reasoning.
 Plant resources (e.g.
-scPlantDB  He et al.
-(2024), Arabidopsis E-CURD-4  (Shulse et al., 2019)) extend modeling beyond animal systems, opening cross-kingdom evaluation.
+scPlantDB  he2024sty, Arabidopsis E-CURD-4  (shulse2019high)) extend modeling beyond animal systems, opening cross-kingdom evaluation.
 Despite this diversity, major gaps persist.
@@
 Foundation models learn transferable cell and gene embeddings directly from large-scale scRNA-seq without explicit labels.
-Representative systems such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) pretrain on multi-tissue atlases exceeding one million cells, using masked-gene or rank-based reconstruction to capture expression context.
-Variants like tGPT  (Shen et al., 2023), scBERT  (Yang et al., 2022), and scGraphformer  (Fan et al., 2024) treat genes as tokens or graph nodes, while cross-species models (UCE  (Rosen et al., 2023), GeneCompass (Yang et al., 2024b)) apply contrastive alignment across homologous genes.
+Representative systems such as scGPT  (cui2024scgpt), Geneformer  (theodoris2023transfer), and scFoundation  (hao2024large) pretrain on multi-tissue atlases exceeding one million cells, using masked-gene or rank-based reconstruction to capture expression context.
+Variants like tGPT  (shen2023generative), scBERT  (yang2022scbert), and scGraphformer  (fan2024scgraphformer) treat genes as tokens or graph nodes, while cross-species models (UCE  (rosen2023universal), GeneCompass (yang2024genecompass)) apply contrastive alignment across homologous genes.
 These models underpin most single-cell LLM pipelines, offering strong transfer for annotation and integration but limited ontology grounding and explainability, motivating later text-bridged and reasoning frameworks.
@@
 Text-Bridge models couple molecular embeddings with biomedical language to ground single-cell representations in semantics and ontology.
-Systems such as scELMo  (Liu et al., 2023), CellLM  (Zhao et al., 2023), and GenePT  (Chen and Zou, 2024) align gene or cell embeddings with textual descriptors, while Cell2Sentence  (Levine et al., 2024) and Cell2Text  (Kharouiche et al., 2025) translate expression profiles into natural-language summaries.
+Systems such as scELMo  (liu2023scelmo), CellLM  (zhao2023large), and GenePT  (chen2024genept) align gene or cell embeddings with textual descriptors, while Cell2Sentence  (pmlr-v235-levine24a) and Cell2Text  (kharouiche2025cell2text) translate expression profiles into natural-language summaries.
 Using dual-encoder or encoder–decoder architectures, these frameworks employ contrastive or prompt-based alignment between molecular and text encoders (e.g., BioBERT).
@@
 Spatial and multimodal frameworks integrate gene expression with spatial coordinates, histology, or additional omics to capture tissue architecture.
-Representative models include TransformerST  (Lu et al., 2024), spaLLM (Ji et al., 2024), and OmiCLIP (Cui et al., 2025) for spatial mapping, and scMMGPT (Shi et al., 2025) and FmH2ST for RNA–ATAC–protein integration.
+Representative models include TransformerST  (lu2024sctrans), spaLLM (10778152), and OmiCLIP (cui2025towards) for spatial mapping, and scMMGPT (shi2025multimodal) and FmH2ST for RNA–ATAC–protein integration.
 They use multi-branch Transformers with modality-specific encoders and cross-attention fusion, trained on datasets such as Visium DLPFC and MERFISH.
@@
 Epigenomic foundation models extend LLM concepts to chromatin-accessibility and regulatory data such as scATAC-seq.
-EpiFoundation  (Wu et al., 2025), EpiBERT (Javed et al., 2025), and EpiAttend (Li et al., 2022) learn cis-regulatory patterns from ENCODE and other compendia, while GeneMamba and scMamba use efficient state-space layers for long-range dependency modeling.
-Cross-omic variants (ChromFound (Jiao et al., 2025), GET (Fu et al., 2025)) jointly embed RNA and ATAC to infer gene-regulatory networks.
+EpiFoundation  (wu2025epifoundation), EpiBERT (javed2025multi), and EpiAttend (li2022epiattend) learn cis-regulatory patterns from ENCODE and other compendia, while GeneMamba and scMamba use efficient state-space layers for long-range dependency modeling.
+Cross-omic variants (ChromFound (jiao2025chromfound), GET (fu2025foundation)) jointly embed RNA and ATAC to infer gene-regulatory networks.
 These models improve biological grounding but remain constrained by sparse data and lack unified benchmarks across regulatory modalities.
@@
 Agentic systems integrate pretrained models with reasoning modules for autonomous single-cell analysis.
-Frameworks such as scAgent (Mao et al., 2025), CellVerse (Zhang et al., 2025b), and Teddy (Chevalier et al., 2025) combine domain-specific encoders with LLM controllers that plan tasks, query ontologies, and interface with tools or APIs.
-EpiAgent (Chen et al., 2025b) extends this paradigm to regulatory genomics.
+Frameworks such as scAgent (mao2025scagent), CellVerse (zhang2025cellverse), and Teddy (chevalier2025teddy) combine domain-specific encoders with LLM controllers that plan tasks, query ontologies, and interface with tools or APIs.
+EpiAgent (chen2025epiagent) extends this paradigm to regulatory genomics.
 These systems enable dialogue-based annotation and multi-step reasoning but lack standardized benchmarks for reasoning fidelity and rely on underlying LLM reliability.
@@
 Foundation models such as scGPT, scFoundation, and scBERT achieve high accuracy on atlases like Tabula Sapiens and the Human Cell Atlas.
-Text-bridge models (CellLM  (Zhao et al., 2023), scELMo  (Liu et al., 2023), GenePT  (Chen and Zou, 2024)) add ontology-based interpretability, and agentic systems (scAgent) perform multi-step annotation reasoning via LLM controllers.
+Text-bridge models (CellLM  (zhao2023large), scELMo  (liu2023scelmo), GenePT  (chen2024genept)) add ontology-based interpretability, and agentic systems (scAgent) perform multi-step annotation reasoning via LLM controllers.
 Models are fine-tuned on curated references (e.g., Azimuth PBMC, HCA Lung) using masked-gene or contrastive objectives.
@@
 Trajectory modeling captures dynamic state transitions and causal responses to interventions.
-Geneformer  (Theodoris et al., 2023), scGPT (Cui et al., 2024), and scRDiT  (Dong et al., 2025) model temporal or perturbation trajectories using denoising and contrastive objectives, trained on datasets such as Replogle 2022 Perturb-seq and sci-Plex.
-Epigenomic models (EpiFoundation (Wu et al., 2025) ,EpiAgent (Chen et al., 2025b)) extend to regulatory responses through chromatin context and ontology reasoning.
+Geneformer  (theodoris2023transfer), scGPT (cui2024scgpt), and scRDiT  (dong2025scrdit) model temporal or perturbation trajectories using denoising and contrastive objectives, trained on datasets such as Replogle 2022 Perturb-seq and sci-Plex.
+Epigenomic models (EpiFoundation (wu2025epifoundation) ,EpiAgent (chen2025epiagent)) extend to regulatory responses through chromatin context and ontology reasoning.
 Metrics include expression-profile correlation and target-gene recovery.
@@
 Multi-omic integration seeks unified representations across paired RNA, ATAC, and protein modalities to capture gene-regulatory and signaling relationships.
-Representative models include scMMGPT (Shi et al., 2025), GET (General Expression Transformer) (Fu et al., 2025), ChromFound (Jiao et al., 2025), and epigenomic-aware variants such as EpiFoundation (Wu et al., 2025) and EpiBERT (Javed et al., 2025).
+Representative models include scMMGPT (shi2025multimodal), GET (General Expression Transformer) (fu2025foundation), ChromFound (jiao2025chromfound), and epigenomic-aware variants such as EpiFoundation (wu2025epifoundation) and EpiBERT (javed2025multi).
 These systems train on joint datasets like 10x Multiome PBMC 10k, TEA-seq, DOGMA-seq, and ASAP-seq, using cross-modal transformers or contrastive alignment between expression and accessibility features.
@@
 Spatial mapping links molecular profiles to their tissue locations, enabling inference of cell organization and microenvironmental context.
-Representative models include TransformerST  (Lu et al., 2024), spaLLM , OmiCLIP, FmH2ST, HEIST  (Madhu et al., 2025), and Spatial2Sentence (Chen et al., 2025a), trained on datasets such as Visium DLPFC, Slide-seqV2, MERFISH, and Xenium.
+Representative models include TransformerST  (lu2024sctrans), spaLLM , OmiCLIP, FmH2ST, HEIST  (madhu2025heist), and Spatial2Sentence (chen2025spatial), trained on datasets such as Visium DLPFC, Slide-seqV2, MERFISH, and Xenium.
 They combine spatial coordinates, histology features, and gene-expression embeddings through cross-attention or contrastive alignment to reconstruct cell or spot-level maps.
@@
 Regulatory-network and pathway inference focuses on identifying gene–gene or enhancer–promoter dependencies that drive transcriptional programs.
-Representative models include GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025), EpiFoundation (Wu et al., 2025), EpiBERT (Javed et al., 2025), ChromFound (Jiao et al., 2025), and GET (Fu et al., 2025), trained on large-scale ATAC and multiome atlases such as ENCODE, human cCRE datasets, and 10x Multiome PBMC.
+Representative models include GeneMamba (qi2025bidirectional), scMamba (yuan2025scmamba), EpiFoundation (wu2025epifoundation), EpiBERT (javed2025multi), ChromFound (jiao2025chromfound), and GET (fu2025foundation), trained on large-scale ATAC and multiome atlases such as ENCODE, human cCRE datasets, and 10x Multiome PBMC.
 These models treat chromatin regions or genes as tokens and learn context-dependent regulatory relationships through self- or cross-attention.
-GeneMamba (Qi et al., 2025a) and scMamba employ state-space layers for efficient modeling of long-range chromatin dependencies, while EpiBERT and EpiFoundation use masked-region reconstruction to capture enhancer–promoter coupling.
+GeneMamba (qi2025bidirectional) and scMamba employ state-space layers for efficient modeling of long-range chromatin dependencies, while EpiBERT and EpiFoundation use masked-region reconstruction to capture enhancer–promoter coupling.
 ChromFound and GET link RNA and accessibility signals to infer transcription-factor activity and causal directionality.
@@
 Cross-species translation aims to transfer cell and gene representations between organisms, enabling comparative and evolutionary analysis of single-cell atlases.
-Representative models include iSEEEK  (Shen et al., 2022), UCE (Universal Cell Embeddings)  (Rosen et al., 2023), GeneCompass  (Yang et al., 2024b), and scPlantLLM  (Cao et al., 2025), trained on paired or homolog-mapped datasets such as tthe Mouse Cell Atlas, Tabula Muris, and scPlantDB.
+Representative models include iSEEEK  (shen2022universal), UCE (Universal Cell Embeddings)  (rosen2023universal), GeneCompass  (yang2024genecompass), and scPlantLLM  (scplantlmm), trained on paired or homolog-mapped datasets such as tthe Mouse Cell Atlas, Tabula Muris, and scPlantDB.
 These systems learn alignment between orthologous genes or conserved cell states to support zero-shot annotation across species.
-iSEEEK  (Shen et al., 2022) integrates over 11 million human and mouse cells using shared gene-token vocabularies, while UCE  (Shen et al., 2023) applies masked autoencoding and contrastive alignment across ortholog groups.
-GeneCompass  (Yang et al., 2024b) extends this with knowledge-informed pretraining that links species via ontology-derived embeddings, and scPlantLLM adapts the paradigm to plant single-cell data across 17 species.
+iSEEEK  (shen2022universal) integrates over 11 million human and mouse cells using shared gene-token vocabularies, while UCE  (shen2023generative) applies masked autoencoding and contrastive alignment across ortholog groups.
+GeneCompass  (yang2024genecompass) extends this with knowledge-informed pretraining that links species via ontology-derived embeddings, and scPlantLLM adapts the paradigm to plant single-cell data across 17 species.
 Evaluation uses transfer accuracy and embedding consistency across species pairs.
@@
 Generative modeling aims to synthesize realistic single-cell profiles, simulate perturbations, or reconstruct missing modalities.
-Representative approaches include scGPT, scFoundation, CellFM  (Zeng et al., 2025), and Geneformer, which employ decoder or diffusion-style architectures to model the probability distribution of gene-expression states.
+Representative approaches include scGPT, scFoundation, CellFM  (zeng2025cellfm), and Geneformer, which employ decoder or diffusion-style architectures to model the probability distribution of gene-expression states.
 These models are pretrained on large atlases such as Tabula Sapiens and HCA, then fine-tuned to generate new cell embeddings or perturbation outcomes.
@@
 Drug-response prediction models aim to infer cellular transcriptional outcomes following chemical or genetic perturbations, supporting virtual screening and mechanism discovery.
-Representative frameworks include Geneformer (Theodoris et al., 2023), scGPT  (Cui et al., 2024), EpiFoundation (Wu et al., 2025), and EpiAgent (Chen et al., 2025b), trained or evaluated on large-scale pharmacotranscriptomic resources such as sci-Plex, Replogle 2022 Perturb-seq, and the ARC Virtual Cell Challenge.
+Representative frameworks include Geneformer (theodoris2023transfer), scGPT  (cui2024scgpt), EpiFoundation (wu2025epifoundation), and EpiAgent (chen2025epiagent), trained or evaluated on large-scale pharmacotranscriptomic resources such as sci-Plex, Replogle 2022 Perturb-seq, and the ARC Virtual Cell Challenge.
 These systems learn conditional embeddings that map control to perturbed states, enabling zero-shot prediction of unseen compounds or targets.
@@
 Robustness to batch effects varies widely.
-scFoundation  (Hao et al., 2024), CellFM  (Zeng et al., 2025), and iSEEEK  (Shen et al., 2022) mitigate donor and protocol variation through large-scale multi-domain pretraining, while others rely on explicit batch tokens or adversarial alignment.
+scFoundation  (hao2024large), CellFM  (zeng2025cellfm), and iSEEEK  (shen2022universal) mitigate donor and protocol variation through large-scale multi-domain pretraining, while others rely on explicit batch tokens or adversarial alignment.
 True cross-platform generalization remains limited; even large models show degradation when tested on unseen sequencing chemistries or tissue types.
 6.3 Multi-omics Alignment
-Models such as scMGPT  (Palayew et al., 2025), GET  (Fu et al., 2025), and ChromFound  (Jiao et al., 2025) demonstrate consistent alignment across RNA, ATAC, and protein features using cross-attention or contrastive fusion.
+Models such as scMGPT  (palayew2025towards), GET  (fu2025foundation), and ChromFound  (jiao2025chromfound) demonstrate consistent alignment across RNA, ATAC, and protein features using cross-attention or contrastive fusion.
 Spatial hybrids (HEIST, FmH2ST) further connect molecular and visual domains.
@@
 Cross-organism transfer remains challenging.
-iSEEEK  (Shen et al., 2022), UCE  (Rosen et al., 2023), and GeneCompass  (Yang et al., 2024b) align human and mouse cell embeddings via ortholog mapping, while scPlantLLM  (Cao et al., 2025) extends this to plants.
+iSEEEK  (shen2022universal), UCE  (rosen2023universal), and GeneCompass  (yang2024genecompass) align human and mouse cell embeddings via ortholog mapping, while scPlantLLM  (scplantlmm) extends this to plants.
 Performance drops sharply for divergent taxa, reflecting incomplete ortholog tables and bias toward human datasets.
@@
 Human and immune-cell–dominant atlases overrepresent certain tissues and demographics.
-Teddy (Chevalier et al., 2025) introduces benchmarking for fairness and representation balance, but fairness metrics are rarely adopted.
+Teddy (chevalier2025teddy) introduces benchmarking for fairness and representation balance, but fairness metrics are rarely adopted.
 No model yet enforces demographic or tissue-specific parity during training.
 6.7 Explainability and Interpretability
-Text-bridge systems (GenePT, CellLM, Cell2Text  (Kharouiche et al., 2025)) improve interpretability by linking embeddings to ontology terms.
+Text-bridge systems (GenePT, CellLM, Cell2Text  (kharouiche2025cell2text)) improve interpretability by linking embeddings to ontology terms.
 Agentic systems (scAgent) extend this via natural-language reasoning and tool explanations.
@@
 Data privacy remains underexplored.
-Synthetic generation (scGFT  (Nouri, 2025)) and open-access chat systems (ChatCell  (Fang et al., 2024)) raise concerns over re-identification and data leakage.
+Synthetic generation (scGFT  (nouri2025single)) and open-access chat systems (ChatCell  (fang2024chatcell)) raise concerns over re-identification and data leakage.
 No single-cell LLM currently implements federated or privacy-preserving learning, and ethical guidance for multi-omic sharing remains informal.
 6.9 Scalability and Efficiency
-Scaling trends mirror NLP: efficient transformers (xTrimoGene (Gong et al., 2023)) and state-space models (GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025)) reduce memory cost while retaining accuracy.
-CellFM  (Zeng et al., 2025) demonstrates training across 100 M cells, showing feasibility of atlas-scale learning.
+Scaling trends mirror NLP: efficient transformers (xTrimoGene (gong2023xtrimogene)) and state-space models (GeneMamba (qi2025bidirectional), scMamba (yuan2025scmamba)) reduce memory cost while retaining accuracy.
+CellFM  (zeng2025cellfm) demonstrates training across 100 M cells, showing feasibility of atlas-scale learning.
 Still, high compute requirements limit accessibility for most research groups.
 6.10 Emerging Paradigms and Agentic Behavior
-Recent frameworks (scAgent  (Mao et al., 2025), CellVerse  (Zhang et al., 2025b), EpiAgent (Chen et al., 2025b)) introduce reasoning and autonomous decision pipelines, integrating LLM controllers with specialized encoders.
+Recent frameworks (scAgent  (mao2025scagent), CellVerse  (zhang2025cellverse), EpiAgent (chen2025epiagent)) introduce reasoning and autonomous decision pipelines, integrating LLM controllers with specialized encoders.
 These systems achieve the highest scores in explainability and cross-modal planning but lack standardized evaluation of reasoning fidelity or reproducibility.
@@
 Future work should establish community leaderboards, causal interpretability tests, and reasoning benchmarks to advance trustworthy and reproducible single-cell intelligence.
+Acknowledgments
+This work was supported in part by Virginia Tech, the Department of Computer Science, and the U.S.
+National Science Foundation (NSF) under Awards #2125798, #2344169, and #2319522.
+We gratefully acknowledge their support and resources that made this research possible.
 Appendix A Data Collection
@@
 tGPT and scBERT encode ranked or discrete gene sequences, while Geneformer and scFoundation use masked-token modeling.
-Hybrid designs such as scGraphformer  (Fan et al., 2024) and scRDiT  (Dong et al., 2025) incorporate graph or diffusion-based attention, improving spatial and regulatory context.
+Hybrid designs such as scGraphformer  (fan2024scgraphformer) and scRDiT  (dong2025scrdit) incorporate graph or diffusion-based attention, improving spatial and regulatory context.
 Learning objectives.
@@
 Agentic models couple (i) a pretrained molecular encoder, (ii) an LLM-based controller, and (iii) external tool interfaces.
-scAgent executes multi-step annotation workflows via ontology queries; CellVerse coordinates multiple domain agents for transcriptomic, spatial, and literature reasoning; EpiAgent extends agentic control to enhancer–gene analysis; and Teddy/scPilot (Joodaki et al., 2024) provide lightweight orchestrators for benchmarking pipelines.
+scAgent executes multi-step annotation workflows via ontology queries; CellVerse coordinates multiple domain agents for transcriptomic, spatial, and literature reasoning; EpiAgent extends agentic control to enhancer–gene analysis; and Teddy/scPilot (joodaki2024detection) provide lightweight orchestrators for benchmarking pipelines.
 Grounding sources include the Human Cell Atlas ontology, UBERON, and PubMed.
@@
 Model (Year) Published where Category Modality Grounding Type Agentic (Y/N) Primary Task Domain Score
-scGPT Cui et al.
-(2024) Nature Methods Foundation Multiomics (scRNA + optionally multi-omic modes) Atlas (trained on large atlas of single-cell datasets) No Annotation, Integration, Perturbation (annotation as primary) 6
-Geneformer Theodoris et al. (2023) Nature Foundation scRNA None / Rank-based (implicit) No Cell classification, in-silico perturbation, network prediction 6
-scFoundation Hao et al.
-(2024) Nature Methods Foundation scRNA / multi-omics (transcriptomics focus) Atlas (large pretrained corpus) No Cell annotation, perturbation prediction, drug response, gene module inference 7
-CellFM Zeng et al. (2025) Nature Communications Foundation scRNA Atlas / value-projection No Learns embeddings from 100M cells and supports annotation, perturbation, gene function 7
-iSEEEK Shen et al.
-(2022) Briefings in Bioinformatics Foundation scRNA Atlas (11.9M cells, human + mouse) No Integrates massive single-cell datasets via gene-ranking similarity to enable scalable cross-dataset embedding and clustering 5
-tGPT Shen et al.
-(2023) iScience Foundation Model (Autoregressive) scRNA Atlas (22.3 M cells) No Treats gene-expression ranks as token sequences and learns generative embeddings for cell clustering, trajectory inference, and bulk-tissue analysis 5
-scBERT Yang et al.
-(2022) Nature Machine Intelligence Foundation Model (Encoder-only) scRNA Atlas (public scRNA-seq corpora) No Learns bidirectional gene–cell embeddings using BERT-style masked modeling for robust cell-type annotation and novel cell discovery 4
-UCE (Universal Cell Embeddings) Rosen et al.
-(2023) BioRxiv Foundation scRNA (cross-species) Atlas / binary masked self-supervision No Embeds any cell zero-shot across species into a shared latent space for clustering, lineage inference, annotation 6
-GeneCompass Yang et al.
-(2024b) Nature-/Cell Research Foundation Model scRNA / cross-species Ontology (GRN + co-expression + gene family knowledge) No Learns cross-species gene–cell embeddings and supports annotation, perturbation, dose-response, and GRN inference 6
-scELMo Liu et al.
-(2023) BioRxiv Text-Bridge LLM scRNA + Metadata Text LLM-generated embeddings from gene/metadata descriptions + raw data No Embeds cells via text-derived gene metadata embeddings combined with expression, then supports clustering, batch correction, annotation, perturbation 5
-CellPLM Wen et al.
-(2023) ICLR Spatial (Multimodal Foundation Model) scRNA + Spatial (SRT) Atlas + Spatial relations (uses SRT in pretraining) No Treats cells as tokens and tissues as sentences, leveraging spatially-resolved transcriptomics and a Gaussian-mixture prior to encode inter-cell relations for denoising, spatial imputation, and perturbation prediction.
+scGPT cui2024scgpt Nature Methods Foundation Multiomics (scRNA + optionally multi-omic modes) Atlas (trained on large atlas of single-cell datasets) No Annotation, Integration, Perturbation (annotation as primary) 6
+Geneformer theodoris2023transfer Nature Foundation scRNA None / Rank-based (implicit) No Cell classification, in-silico perturbation, network prediction 6
+scFoundation hao2024large Nature Methods Foundation scRNA / multi-omics (transcriptomics focus) Atlas (large pretrained corpus) No Cell annotation, perturbation prediction, drug response, gene module inference 7
+CellFM zeng2025cellfm Nature Communications Foundation scRNA Atlas / value-projection No Learns embeddings from 100M cells and supports annotation, perturbation, gene function 7
+iSEEEK shen2022universal Briefings in Bioinformatics Foundation scRNA Atlas (11.9M cells, human + mouse) No Integrates massive single-cell datasets via gene-ranking similarity to enable scalable cross-dataset embedding and clustering 5
+tGPT shen2023generative iScience Foundation Model (Autoregressive) scRNA Atlas (22.3 M cells) No Treats gene-expression ranks as token sequences and learns generative embeddings for cell clustering, trajectory inference, and bulk-tissue analysis 5
+scBERT yang2022scbert Nature Machine Intelligence Foundation Model (Encoder-only) scRNA Atlas (public scRNA-seq corpora) No Learns bidirectional gene–cell embeddings using BERT-style masked modeling for robust cell-type annotation and novel cell discovery 4
+UCE (Universal Cell Embeddings) rosen2023universal BioRxiv Foundation scRNA (cross-species) Atlas / binary masked self-supervision No Embeds any cell zero-shot across species into a shared latent space for clustering, lineage inference, annotation 6
+GeneCompass yang2024genecompass Nature-/Cell Research Foundation Model scRNA / cross-species Ontology (GRN + co-expression + gene family knowledge) No Learns cross-species gene–cell embeddings and supports annotation, perturbation, dose-response, and GRN inference 6
+scELMo liu2023scelmo BioRxiv Text-Bridge LLM scRNA + Metadata Text LLM-generated embeddings from gene/metadata descriptions + raw data No Embeds cells via text-derived gene metadata embeddings combined with expression, then supports clustering, batch correction, annotation, perturbation 5
+CellPLM wen2023cellplm ICLR Spatial (Multimodal Foundation Model) scRNA + Spatial (SRT) Atlas + Spatial relations (uses SRT in pretraining) No Treats cells as tokens and tissues as sentences, leveraging spatially-resolved transcriptomics and a Gaussian-mixture prior to encode inter-cell relations for denoising, spatial imputation, and perturbation prediction.
 4
-scMoFormer Tang et al.
-(2023) ArXiv Spatial (Multimodal Foundation Model) scRNA + Protein / Multi-omics Atlas + domain knowledge in cross-modality aggregation No Uses modality-specific transformers and cross-attention to impute missing modalities, classify cells, and fuse multimodal representations 4
-scFormer Xu et al.
-(2024) ArXiv Spatial (Multimodal Foundation Model) (or Foundation + multi-omics) Transcript- omics + Proteomics / multimodal Atlas + transformer-based fusion No Aligns and integrates multi-omics single-cell data, recovers missing modalities, and transfers labels across modalities 4
-scMulan Bian et al.
-(2024a) BioRxiv Foundation Model scRNA + Metadata Atlas + prompt-conditioned generative modeling No Encodes each cell as a “c-sentence” integrating expression + metadata; supports zero-shot annotation, batch integration, and conditional generation 3
-scPRINT Kalfon et al.
-(2025) Nature Communications Foundation Model scRNA Atlas (50M+ cell pretraining) No Learns cell embeddings and infers cell-specific gene regulatory networks; supports zero-shot denoising, batch correction, label prediction, expression reconstruction 5
-scGraphformer Fan et al.
-(2024) Nature Communications Biology Foundation Model scRNA Atlas + relational prior (kNN bias, refined) No Learns a cell–cell graph via a transformer-GNN hybrid for better classification and interaction inference 4
-scRDiT Dong et al.
-(2025) ArXiv Foundation Model scRNA (transcriptome) Atlas-like generative prior / diffusion modeling No Generates synthetic scRNA-seq samples via diffusion transformer + DDIM for accelerated sampling 2
-scGFT Nouri (2025) Nature Communications Biology Foundation Model scRNA Fourier-based perturbation / reconstruction (train-free) No Synthesizes new single-cell expression profiles by perturbing Fourier components in frequency space.
+scMoFormer tang2023single ArXiv Spatial (Multimodal Foundation Model) scRNA + Protein / Multi-omics Atlas + domain knowledge in cross-modality aggregation No Uses modality-specific transformers and cross-attention to impute missing modalities, classify cells, and fuse multimodal representations 4
+scFormer xu2024scmformer ArXiv Spatial (Multimodal Foundation Model) (or Foundation + multi-omics) Transcript- omics + Proteomics / multimodal Atlas + transformer-based fusion No Aligns and integrates multi-omics single-cell data, recovers missing modalities, and transfers labels across modalities 4
+scMulan bian2024scmulan BioRxiv Foundation Model scRNA + Metadata Atlas + prompt-conditioned generative modeling No Encodes each cell as a “c-sentence” integrating expression + metadata; supports zero-shot annotation, batch integration, and conditional generation 3
+scPRINT kalfon2025scprint Nature Communications Foundation Model scRNA Atlas (50M+ cell pretraining) No Learns cell embeddings and infers cell-specific gene regulatory networks; supports zero-shot denoising, batch correction, label prediction, expression reconstruction 5
+scGraphformer fan2024scgraphformer Nature Communications Biology Foundation Model scRNA Atlas + relational prior (kNN bias, refined) No Learns a cell–cell graph via a transformer-GNN hybrid for better classification and interaction inference 4
+scRDiT dong2025scrdit ArXiv Foundation Model scRNA (transcriptome) Atlas-like generative prior / diffusion modeling No Generates synthetic scRNA-seq samples via diffusion transformer + DDIM for accelerated sampling 2
+scGFT nouri2025single Nature Communications Biology Foundation Model scRNA Fourier-based perturbation / reconstruction (train-free) No Synthesizes new single-cell expression profiles by perturbing Fourier components in frequency space.
 4
-scTrans Lu et al.
-(2024) IJCAI Foundation Model scRNA Sub-vector masked completion over gene modules No Learns multi-scale sub-vector tokens to perform gene-selective cell-type annotation via masked completion and contrastive regularization 3
-scGT (Graph Transformer) Qi et al. (2025b) Bioinfo-rmatics Advance Graph / Multi-omics Integration scRNA + scATAC Observed / Hybrid Graph No Multi-omics integration + label transfer 4
-TransformerST Lu et al.
-(2024) Briefings in Bioinformatics Spatial / Multimodal FM Spatial (histology + gene expression) Spatial (histology + gene expression) No Super-resolution gene expression & tissue clustering 3
-scGPT-spatial Cui et al.
-(2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial Transcriptomics + scRNA prior Atlas + spatial-aware decoding No Extends scGPT via continual pretraining to spatial data, supports multi-slide integration, cell-type deconvolution, and spatial gene imputation 3
-HEIST Madhu et al.
-(2025) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + proteomics Hierarchical graph modeling (spatial + GRNs) No Learns joint embeddings of cells and genes in spatial context to perform cell annotation, gene imputation, spatial clustering, clinical outcome prediction 4
-stFormer Cao et al.
-(2024) BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + ligand context Atlas + biased cross-attention to ligand niche genes No Learns gene embeddings contextualized by ligand signals in spatial microenvironments; aids clustering, ligand–receptor inference, and perturbation simulation 4
-FmH2ST Wang et al.
-(2025) Nucleic Acids Research Spatial (Multimodal Foundation Model) Histology image + spatial transcriptomics Image foundation + dual graphs + spot branch No Predict spatial gene expression from histology using fused image and spot features, supporting denoising, heterogeneity detection, and regulatory inference 3
-OmiCLIP Cui et al.
-(2025) Nature Methods Spatial (Multimodal Foundation Model) Histology + Spatial Transcriptomics Image–gene contrastive alignment (rank-based) No Learns unified visual-omics embeddings linking histopathology and spatial gene expression; enables cross-modal prediction, annotation, and tissue retrieval 7
-QuST-LLM Huang (2024) ArXiv Text-Bridge LLM Spatial transcriptomics + histology metadata GO-term + gene enrichment + LLM narrative overlay Yes Converts ST data and ROIs into human-readable narratives and matches natural language queries to spatial regions via GO/LLM interpretation 4
-SpaCCC Yang et al.
-(2024a) IEEE Xplore Text-Bridge LLM Spatial transcriptomics / transcriptomic genes (LRs) LLM embeddings of ligand + receptor genes No Infers spatially resolved cell–cell communication by embedding LR pairs in LLM latent space + diffusion / permutation test filtering 3
-spaLLM Ji et al.
-(2024) Briefings in Bioinformatics Spatial (Multimodal Foundation Model) Spatial multi-omics (RNA, ATAC, protein) scGPT embeddings + GNN + attention fusion No Enhances spatial domain identification by fusing LLM-derived embeddings and spatial-omics signals via multi-view attention 3
-scMMGPT (2025) Shi et al.
-(2025) ArXiv Text-Bridge LLM / Multimodal_FM RNA + Text (metadata) Annotation, description No Generation - links single-cell and text PLMs to describe cells, generate pseudo-cells from text, and enhance annotation through text-conditioned reasoning 6
-Cell2Sentence (C2S) Levine et al.
-(2024) PMLR Text-Bridge LLM RNA (scRNA-seq converted to “cell sentences”) Marker / Implicit (gene rank-order) No Generation, annotation & reconstruction - encodes cells as gene-ranked sentences and fine-tunes LLMs to classify or generate biologically meaningful cell text 6
-C2S-Scale Rizvi et al.
-(2025) BioRxiv Text-Bridge LLM+ Multimodal Foundation Model RNA + Text / Metadata Atlas + Text / Implicit rank grounding No Chat, generation & annotation - trains large LLMs on cell sentences and biological text to enable perturbation prediction and multicellular summarization 7
-Cell2Text Kharouiche et al.
-(2025) ArXiv Text-Bridge LLM + Multimodal Foundation Model RNA → Natural Language Gene-level embeddings + ontology / metadata grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict expression and interpret TF–cis interactions 7
-GenePT Chen and Zou (2024) BioRxiv Foundation Model / Text-Augmented scRNA Literature / Text embedding grounding No Embedding & downstream prediction - uses GPT-3.5 gene text embeddings to derive cell embeddings via weighted or ranked gene aggregation.
+scTrans lu2024sctrans IJCAI Foundation Model scRNA Sub-vector masked completion over gene modules No Learns multi-scale sub-vector tokens to perform gene-selective cell-type annotation via masked completion and contrastive regularization 3
+scGT (Graph Transformer) qi2025scgt Bioinfo-rmatics Advance Graph / Multi-omics Integration scRNA + scATAC Observed / Hybrid Graph No Multi-omics integration + label transfer 4
+TransformerST lu2024sctrans Briefings in Bioinformatics Spatial / Multimodal FM Spatial (histology + gene expression) Spatial (histology + gene expression) No Super-resolution gene expression & tissue clustering 3
+scGPT-spatial cui2024scgpt BioRxiv Spatial (Multimodal Foundation Model) Spatial Transcriptomics + scRNA prior Atlas + spatial-aware decoding No Extends scGPT via continual pretraining to spatial data, supports multi-slide integration, cell-type deconvolution, and spatial gene imputation 3
+HEIST madhu2025heist BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + proteomics Hierarchical graph modeling (spatial + GRNs) No Learns joint embeddings of cells and genes in spatial context to perform cell annotation, gene imputation, spatial clustering, clinical outcome prediction 4
+stFormer cao2024stformer BioRxiv Spatial (Multimodal Foundation Model) Spatial transcriptomics + ligand context Atlas + biased cross-attention to ligand niche genes No Learns gene embeddings contextualized by ligand signals in spatial microenvironments; aids clustering, ligand–receptor inference, and perturbation simulation 4
+FmH2ST wang2025fmh2st Nucleic Acids Research Spatial (Multimodal Foundation Model) Histology image + spatial transcriptomics Image foundation + dual graphs + spot branch No Predict spatial gene expression from histology using fused image and spot features, supporting denoising, heterogeneity detection, and regulatory inference 3
+OmiCLIP cui2025towards Nature Methods Spatial (Multimodal Foundation Model) Histology + Spatial Transcriptomics Image–gene contrastive alignment (rank-based) No Learns unified visual-omics embeddings linking histopathology and spatial gene expression; enables cross-modal prediction, annotation, and tissue retrieval 7
+QuST-LLM huang2024qust ArXiv Text-Bridge LLM Spatial transcriptomics + histology metadata GO-term + gene enrichment + LLM narrative overlay Yes Converts ST data and ROIs into human-readable narratives and matches natural language queries to spatial regions via GO/LLM interpretation 4
+SpaCCC yang2024deciphering IEEE Xplore Text-Bridge LLM Spatial transcriptomics / transcriptomic genes (LRs) LLM embeddings of ligand + receptor genes No Infers spatially resolved cell–cell communication by embedding LR pairs in LLM latent space + diffusion / permutation test filtering 3
+spaLLM 10778152 Briefings in Bioinformatics Spatial (Multimodal Foundation Model) Spatial multi-omics (RNA, ATAC, protein) scGPT embeddings + GNN + attention fusion No Enhances spatial domain identification by fusing LLM-derived embeddings and spatial-omics signals via multi-view attention 3
+scMMGPT (2025) shi2025multimodal ArXiv Text-Bridge LLM / Multimodal_FM RNA + Text (metadata) Annotation, description No Generation - links single-cell and text PLMs to describe cells, generate pseudo-cells from text, and enhance annotation through text-conditioned reasoning 6
+Cell2Sentence (C2S) pmlr-v235-levine24a PMLR Text-Bridge LLM RNA (scRNA-seq converted to “cell sentences”) Marker / Implicit (gene rank-order) No Generation, annotation & reconstruction - encodes cells as gene-ranked sentences and fine-tunes LLMs to classify or generate biologically meaningful cell text 6
+C2S-Scale Rizvi2025.04.14.648850 BioRxiv Text-Bridge LLM+ Multimodal Foundation Model RNA + Text / Metadata Atlas + Text / Implicit rank grounding No Chat, generation & annotation - trains large LLMs on cell sentences and biological text to enable perturbation prediction and multicellular summarization 7
+Cell2Text kharouiche2025cell2text ArXiv Text-Bridge LLM + Multimodal Foundation Model RNA → Natural Language Gene-level embeddings + ontology / metadata grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict expression and interpret TF–cis interactions 7
+GenePT chen2024genept BioRxiv Foundation Model / Text-Augmented scRNA Literature / Text embedding grounding No Embedding & downstream prediction - uses GPT-3.5 gene text embeddings to derive cell embeddings via weighted or ranked gene aggregation.
 5-7
-CellLM Zhao et al.
-(2023) ArXiv Foundation Model / Representation model RNA (single cell expression) Implicit - contrastive embedding from expression data No Represent cells via a contrastive-learning transformer, optimize embedding space for tasks like cell-type annotation, drug sensitivity prediction 6
-scExtract Wu and Tang (2025) Genome Biology Agentic Framework (Text-Bridge LLM) RNA (scRNA-seq) + Text Article-based parameter extraction Yes Annotation & Integration - uses LLMs to extract pipeline parameters from publications and apply them for dataset harmonization.
+CellLM zhao2023large ArXiv Foundation Model / Representation model RNA (single cell expression) Implicit - contrastive embedding from expression data No Represent cells via a contrastive-learning transformer, optimize embedding space for tasks like cell-type annotation, drug sensitivity prediction 6
+scExtract wu2025scextract Genome Biology Agentic Framework (Text-Bridge LLM) RNA (scRNA-seq) + Text Article-based parameter extraction Yes Annotation & Integration - uses LLMs to extract pipeline parameters from publications and apply them for dataset harmonization.
 7
-scAgent Mao et al.
-(2025) ArXiv Agentic Framework (Text-Bridge hybrid) RNA (scRNA-seq) Reference atlas + marker gene reasoning + memory grounding Yes Universal cell annotation & novel cell discovery - scAgent uses an LLM planning module, memory, and tool modules to annotate cells across tissues, detect unknown types, and incrementally learn new annotations 7
-EpiFoundation Wu et al.
-(2025) BioRxiv Epigenomic Foundation Model / Multimodal alignment scATAC (chromatin accessibility) Peak-to-gene supervision (alignment to expression) No Cell embedding, annotation, batch correction, gene expression prediction - trains on sparse peak sets, aligns to gene expression supervision, and transfers learned embeddings for downstream ATAC tasks 7
-EpiAgent Chen et al.
-(2025b) BioRxiv Epigenomic Foundation Model / Multimodal text-bridge hybrid scATAC / chromatin accessibility Peak tokenization + external embeddings + regulatory supervision grounding No Embedding, annotation, imputation, and perturbation prediction - encodes chromatin accessibility as ranked cCRE tokens, enabling zero-shot cell annotation, peak imputation, and response prediction.
+scAgent mao2025scagent ArXiv Agentic Framework (Text-Bridge hybrid) RNA (scRNA-seq) Reference atlas + marker gene reasoning + memory grounding Yes Universal cell annotation & novel cell discovery - scAgent uses an LLM planning module, memory, and tool modules to annotate cells across tissues, detect unknown types, and incrementally learn new annotations 7
+EpiFoundation wu2025epifoundation BioRxiv Epigenomic Foundation Model / Multimodal alignment scATAC (chromatin accessibility) Peak-to-gene supervision (alignment to expression) No Cell embedding, annotation, batch correction, gene expression prediction - trains on sparse peak sets, aligns to gene expression supervision, and transfers learned embeddings for downstream ATAC tasks 7
+EpiAgent chen2025epiagent BioRxiv Epigenomic Foundation Model / Multimodal text-bridge hybrid scATAC / chromatin accessibility Peak tokenization + external embeddings + regulatory supervision grounding No Embedding, annotation, imputation, and perturbation prediction - encodes chromatin accessibility as ranked cCRE tokens, enabling zero-shot cell annotation, peak imputation, and response prediction.
 8
-ChromFound Jiao et al.
-(2025) Nature Epigenomic → Expression Transformer / Regulatory FM scATAC (chromatin accessibility) + DNA sequence Motif × peak matrix / masked regulatory grammar grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict gene expression in seen and unseen cell types; also interprets transcription factor interactions and cis-regulatory elements 8
-GET (General Expression Transformer) Fu et al.
-(2025) Nature Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
-scMPT Palayew et al.
-(2025) ArXiv Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
-EpiBERT Javed et al.
-(2025) Cell Genomics Epigenomic Foundation Model / Multi-modal Transformer DNA sequence + chromatin accessibility Masked-accessibility pretraining + motif & sequence fusion No Accessibility imputation, gene expression prediction & regulatory inference - predicts masked ATAC signals, then fine-tunes to predict expression and enhancer-gene links, generalizing to unseen cell types 8
-scMamba Yuan et al.
-(2025) ArXiv Multimodal / Foundation_FM Multi-omics (RNA + others) Implicit via integrated features; no explicit external grounding No Multi-omic embedding, integration & annotation across modalities without prior feature selection 7
-GeneMamba Qi et al.
-(2025a) ArXiv Foundation Model / State-Space Model RNA (scRNA) Implicit (via gene sequence context + pathway loss) No Multi-batch integration, cell-type annotation, gene correlation - uses a BiMamba state-space architecture with pathway-aware losses for scalable, context-rich modeling of scRNA 7
-Nicheformer Schaar et al.
-(2024) BioRxiv Foundation _FM/ Spatial / Multimodal_FM scRNA + spatial transcriptomics Contextual tokenization + metadata + spatial neighborhood embedding No Spatial context prediction, spatial label / niche prediction, mapping spatial info to dissociated cells 7
-scFormer Cell+ Cui et al.
-(2022) Bioarxiv Foundation Model / Multimodal Transformer RNA (+ metadata) Joint gene-cell embedding with metadata tokens No Integration & Annotation - context-aware joint embedding for cross-species/tissue generalization 7
-scPlantLLM Cao et al.
-(2025) Genomics, Proteomics and Bioinformatics Text-Bridge LLM (Foundation Model) RNA (plant scRNA-seq) Gene token + binned expression embedding (Plant Atlas grounding) No Annotation, Integration, GRN inference - pretrains on plant single-cell data with masked LM and cell-type supervision, enabling cross-species annotation, clustering, and regulatory discovery in plant systems 7
-LICT Ye et al.
-(2024) ArXiv Text-Bridge LLM / Annotation LLM Hybrid RNA (scRNA-seq) Marker gene–based DE lists + LLM prompting No Annotation & Reliability - iterative LLM prompting with DE markers for label refinement and confidence scoring 5
-Teddy (family of models) Chevalier et al.
-(2025) ArXiv Foundation Model / Disease-aware Transformer scRNA (single-cell RNA-seq) Self-supervised + supervised annotation supervision No Disease state classification / healthy vs diseased detection - trained to identify disease conditions of held-out donors and distinguish diseased vs healthy cells in new disease contexts 7
-Pilot Joodaki et al.
-(2024) Github Benchmark / Evaluation Framework RNA Model-agnostic pilot foundation testing No Benchmarking & Evaluation - lightweight pilot framework for early single-cell FM testing 4
-CellVerse Zhang et al. (2025b) ArXiv Benchmark / Evaluation Multi-omics (RNA, CITE, ASAP, etc.) Implicit via prompt encoding No QA benchmark for annotation, drug response, perturbation tasks 5
-xTrimoGene Gong et al.
-(2023) NeurIPS / ArXiv Foundation Model (Scalable Transformer) scRNA-seq Sparse masking + auto-discretization No Representation learning + annotation, perturbation, drug synergy prediction 7
-EpiAttend Li et al.
-(2022) NeurIPS 2022 Workshop Regulatory / Sequence-Epigenome Transformer DNA sequence + single-cell epigenomic data Sequence + cell-specific epigenome grounding No Predict cell type–specific gene expression by integrating DNA sequence and single-cell epigenomic tracks, linking enhancers and promoters 6
-Spatial2-Sentence Chen et al.
-(2025a) ArXiv Spatial / Text-Bridge hybrid Spatial + expression (Imaging Mass Cytometry) Spatial adjacency + expression similarity tokenization No Encode spatial & expression context into multi-sentence prompts for LLMs to perform cell-type classification and clinical status prediction 6
-ChatCell Fang et al.
-(2024) ArXiv Text-Bridge LLM (Instructional) scRNA → “cell sentence” Vocabulary adaptation + unified sequence generation of cell sentences No Natural language interface for single-cell tasks - allows users to query, annotate, generate, and explore scRNA data via text prompts.
+ChromFound jiao2025chromfound Nature Epigenomic → Expression Transformer / Regulatory FM scATAC (chromatin accessibility) + DNA sequence Motif × peak matrix / masked regulatory grammar grounding No Expression prediction & regulatory inference - learns regulatory syntax from chromatin accessibility and DNA motifs to predict gene expression in seen and unseen cell types; also interprets transcription factor interactions and cis-regulatory elements 8
+GET (General Expression Transformer) fu2025foundation Nature Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
+scMPT palayew2025towards ArXiv Text-Bridge LLM (Hybrid Fusion Model) RNA (scRNA) / Text embeddings Fusion of scGPT embeddings + text-encoded “cell sentences” grounding No Annotation (fusion embedding model) - combines scGPT-derived embeddings and LLM (text encoder) embeddings via a small fusion MLP to improve cell-type classification robustness across datasets 6
+EpiBERT javed2025multi Cell Genomics Epigenomic Foundation Model / Multi-modal Transformer DNA sequence + chromatin accessibility Masked-accessibility pretraining + motif & sequence fusion No Accessibility imputation, gene expression prediction & regulatory inference - predicts masked ATAC signals, then fine-tunes to predict expression and enhancer-gene links, generalizing to unseen cell types 8
+scMamba yuan2025scmamba ArXiv Multimodal / Foundation_FM Multi-omics (RNA + others) Implicit via integrated features; no explicit external grounding No Multi-omic embedding, integration & annotation across modalities without prior feature selection 7
+GeneMamba qi2025bidirectional ArXiv Foundation Model / State-Space Model RNA (scRNA) Implicit (via gene sequence context + pathway loss) No Multi-batch integration, cell-type annotation, gene correlation - uses a BiMamba state-space architecture with pathway-aware losses for scalable, context-rich modeling of scRNA 7
+Nicheformer schaar2024nicheformer BioRxiv Foundation _FM/ Spatial / Multimodal_FM scRNA + spatial transcriptomics Contextual tokenization + metadata + spatial neighborhood embedding No Spatial context prediction, spatial label / niche prediction, mapping spatial info to dissociated cells 7
+scFormer Cell+ cui2022scformer Bioarxiv Foundation Model / Multimodal Transformer RNA (+ metadata) Joint gene-cell embedding with metadata tokens No Integration & Annotation - context-aware joint embedding for cross-species/tissue generalization 7
+scPlantLLM scplantlmm Genomics, Proteomics and Bioinformatics Text-Bridge LLM (Foundation Model) RNA (plant scRNA-seq) Gene token + binned expression embedding (Plant Atlas grounding) No Annotation, Integration, GRN inference - pretrains on plant single-cell data with masked LM and cell-type supervision, enabling cross-species annotation, clustering, and regulatory discovery in plant systems 7
+LICT ye2024objectively ArXiv Text-Bridge LLM / Annotation LLM Hybrid RNA (scRNA-seq) Marker gene–based DE lists + LLM prompting No Annotation & Reliability - iterative LLM prompting with DE markers for label refinement and confidence scoring 5
+Teddy (family of models) chevalier2025teddy ArXiv Foundation Model / Disease-aware Transformer scRNA (single-cell RNA-seq) Self-supervised + supervised annotation supervision No Disease state classification / healthy vs diseased detection - trained to identify disease conditions of held-out donors and distinguish diseased vs healthy cells in new disease contexts 7
+Pilot joodaki2024detection Github Benchmark / Evaluation Framework RNA Model-agnostic pilot foundation testing No Benchmarking & Evaluation - lightweight pilot framework for early single-cell FM testing 4
+CellVerse zhang2025cellverse ArXiv Benchmark / Evaluation Multi-omics (RNA, CITE, ASAP, etc.) Implicit via prompt encoding No QA benchmark for annotation, drug response, perturbation tasks 5
+xTrimoGene gong2023xtrimogene NeurIPS / ArXiv Foundation Model (Scalable Transformer) scRNA-seq Sparse masking + auto-discretization No Representation learning + annotation, perturbation, drug synergy prediction 7
+EpiAttend li2022epiattend NeurIPS 2022 Workshop Regulatory / Sequence-Epigenome Transformer DNA sequence + single-cell epigenomic data Sequence + cell-specific epigenome grounding No Predict cell type–specific gene expression by integrating DNA sequence and single-cell epigenomic tracks, linking enhancers and promoters 6
+Spatial2-Sentence chen2025spatial ArXiv Spatial / Text-Bridge hybrid Spatial + expression (Imaging Mass Cytometry) Spatial adjacency + expression similarity tokenization No Encode spatial & expression context into multi-sentence prompts for LLMs to perform cell-type classification and clinical status prediction 6
+ChatCell fang2024chatcell ArXiv Text-Bridge LLM (Instructional) scRNA → “cell sentence” Vocabulary adaptation + unified sequence generation of cell sentences No Natural language interface for single-cell tasks - allows users to query, annotate, generate, and explore scRNA data via text prompts.
 Hugging Face 6
-CellAtria Nouri et al. (2025) BioRxiv Agentic Framework scRNA- seq + metadata Ontology (graph + metadata) Yes Annotation & Ontology Mapping 7
-CellAgent Xiao et al. (2024) ArXiv Agentic Framework scRNA- seq None Yes Annotation & Ontology Mapping 7
+CellAtria nouri2025agentic BioRxiv Agentic Framework scRNA- seq + metadata Ontology (graph + metadata) Yes Annotation & Ontology Mapping 7
+CellAgent xiao2024cellagent ArXiv Agentic Framework scRNA- seq None Yes Annotation & Ontology Mapping 7
 Appendix F Appendix Figure